### v1

# on the Radon-Nikodym Derivative

Abstract
Rigorous statements and formal proofs are presented for both foundational and advanced folklore theorems on the Radon-Nikodym derivative. The cases of product and marginal measures are carefully considered; and the hypothesis under which the statements hold are rigorously enumerated.
I Introduction
In mathematics, folklore theorems refer to results that are widely accepted and frequently utilized by experts but are often not formally proven or explicitly documented. In game theory for example, the original Folk theorem earned its name because, although it was widely recognized among game theorists during the 1950s, it remains unpublished and without attribution to particular authors. See for instance [1] and [2]. Folklore theorems populate all areas in mathematics. In information theory, a large set of folklore theorems involve the Radon-Nikodym derivative (RND), first introduced by Radon [3]; and later generalized by Nikodym [4]. The existence of such folklore theorems in this area arises in part from the fact that all Shannon’s information measures can be defined in terms of the RND. Interestingly, most properties of the RND are often presented as comments in most textbooks on measure theory and probability theory, c.f, [5, 6, 7, 8, 9, 10, 11]
Claude Shannon did not use the RND in his foundational papers [12, 13] to define entropy and mutual information. Instead, Shannon opted for restricting his publications to the case in which measures either possess a probability mass function or a probability density function, which are both instances of RNDs. This choice significantly influenced the presentation of most subsequent results in information theory and established the style in which classical textbooks were written [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30]. Nonetheless, the RND has been increasingly adopted in modern textbooks [31] and in the definition of new information measures, e.g., lautum information [32], to privilege a unified presentation. That is, independently of the measure used as a reference, e.g., the counting measure, the Lesbegue measure, etc. Some recent results whose presentation relies on the RND are for instance [33, 34, 35, 36, 37, 38, 39, 40, 41, 42]. In the light of the central role of the RND in information theory, this paper presents rigorous statements and formal proofs for the most common folklore theorems on the RND. These include the change of measure theorem, the proportional-measures theorem, the chain rule, the multiplicative inverse, linearity and continuity theorems, as well as the product-measure theorem. Less common theorems such as the unit measure and two Bayes-like rules of the RND are also formally proved. This is the first time such material appears in peer-reviewed literature. Finally, it is important to highlight that particular attention has been put on rigorously stating the conditions under which these theorems hold, hopefully providing a valuable reference for researchers and students in information theory.
II Preliminaries
This section introduces relevant notational conventions alongside the Radon-Nikodym theorem. In particular, some equalities presented in this paper are valid almost surely with respect to a given measure. For clarity, given a measure space $\left(\Omega,\mathscr{F},P\right)$, the notation =$\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{P}$ is introduced and shall be read as “equal for all $x\in\Omega$ except on a negligible set with respect to $P$”; or equivalently as “equal almost surely with respect to $P$”. Moreover, given two measures $P$ and $Q$ on the same measurable space, the notation ${P}\ll{Q}$ stands for “the measure $P$ is absolutely continuous with respect to $Q$”. Using this notation, the Radon-Nikodym derivative is introduced by the following theorem.
Theorem 1 (Radon-Nikodym theorem, [5, Theorem 2.2.1])
Let $P$ and $Q$ be two measures on a given measurable space $\left(\Omega,\mathscr{F}\right)$, such that $Q$ is $\sigma$-finite and ${P}\ll{Q}$. Then, there exists a nonnegative Borel measurable function $g:~{}\Omega~{}\to~{}\mathbb{R}$ such that for all $\mathcal{A}\in\mathscr{F}$,
$P(\mathcal{A})=\int_{\mathcal{A}}g(x)\mathrm{d}Q(x).$ (1)
Moreover, if another function $h$ satisfies for all $\mathcal{A}\in\mathscr{F}$ that $P(\mathcal{A})=\int_{\mathcal{A}}h(x)\mathrm{d}Q(x)$, then $g(x)\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{Q}$$h(x)$.
The function $g$ in (1) is often referred to as the Radom-Nikodym derivative of $P$ with respect to $Q$; and is also written as $\frac{\mathrm{d}P}{\mathrm{d}Q}$, such that $g(x)~{}=~{}\frac{\mathrm{d}P}{\mathrm{d}Q}(x).$
The Radon-Nikodym theorem is the foundational tool from which many folklore theorems in information theory originate. Some of these folklore theorems are thoroughly studied in the following sections.
III Basic Folklore Theorems
This section focuses on basic folklore theorems, where “basic” denotes their well-established nature. One of the most common folklore theorems is often referred to as the “change of measure” theorem.
Theorem 2 (Change of Measure)
Let $P$ and $Q$ be two measures on the measurable space $\left(\Omega,\mathscr{F}\right)$ with ${P}\ll{Q}$; and $Q$ a $\sigma$-finite measure. Let $f:\Omega\to\mathds{R}$ be a Borel measurable function such that the integral $\int_{\Omega}f(x)\mathrm{d}P(x)$ exists. Then, for all $\mathcal{A}\in\mathscr{F}$,
$\int_{\mathcal{A}}f(x)\mathrm{d}P(x)=\displaystyle\int_{\mathcal{A}}f(x)\frac{\mathrm{d}P}{\mathrm{d}Q}(x)\mathrm{d}Q(x).$ (2)
Proof:
[…]
$\int_{\mathcal{A}}f(x)\frac{\mathrm{d}P}{\mathrm{d}Q}(x)\mathrm{d}Q(x)$ $=$ $\int_{\mathcal{A}}\frac{\mathrm{d}P}{\mathrm{d}Q}(x)\sum_{i=1}^{n}a_{i}\mathds{1}_{\mathcal{A}_{i}}(x)\mathrm{d}Q(x)$ (3)
$=$ $\sum_{i=1}^{n}a_{i}\int_{\mathcal{B}_{i}}\frac{\mathrm{d}P}{\mathrm{d}Q}(x)\mathrm{d}Q(x)$ (4)
$=$ $\sum_{i=1}^{n}a_{i}P(\mathcal{B}_{i}),$ (5)
where the equality in (LABEL:EqMartinezPabon102) follows from the linearity of the integral [5, Theorem 1.6.3]; and the equality in (LABEL:EqMartinezPabon105) follows from Theorem 1. On the other hand, for all $\mathcal{A}\in\mathscr{F}$
$\displaystyle\int_{\mathcal{A}}f(x)\mathrm{d}P(x)$ $=$ $\displaystyle\int_{\mathcal{A}}\displaystyle\sum_{i=1}^{n}a_{i}\mathds{1}_{\mathcal{A}_{i}(x)}\mathrm{d}P(x)$ (6)
$=$ $\displaystyle\sum_{i=1}^{n}\int_{\mathcal{A}}a_{i}\mathds{1}_{\mathcal{A}_{i}(x)}\mathrm{d}P(x)$ (7)
$=$ $\sum_{i=1}^{n}a_{i}\int_{\mathcal{B}_{i}}\mathrm{d}P(x)=\sum_{i=1}^{n}a_{i}P(\mathcal{B}_{i}),$ (8)
where the equality in (7) follows from the linearity of the integral [5, Theorem 1.6.3]. Hence, from Theorem 1, and equalities (LABEL:EqMartinezPabon105) and (8), it follows that when $f$ is a simple function, the equality in (2) holds. This concludes the first part of the proof.
The second part of the proof proceeds by considering the following observations: $\left(a\right)$ simple functions form a dense subset of the space of Borel measurable functions [5, Theorem 1.5.5(b)]; and $\left(b\right)$ the integral is a continuous map from that space [5, Theorem 1.6.2]. Hence, from $\left(a\right)$ and $\left(b\right)$, it follows that (2) also holds for any Borel measurable function $f$. This completes the proof. ∎
Another reputed folklore theorem, which is often referred to as the “proportional measures” theorem, establishes the explicit forms of the Radon-Nikodym derivatives between two measures, in which one is proportional to the other.
Theorem 3 (Proportional Measures)
Let $P$ and $Q$ be two $\sigma$- finite measures on the measurable space $\left(\Omega,\mathscr{F}\right)$, such that for all $\mathcal{A}\in\mathscr{F}$
$Q(\mathcal{A})=cP(\mathcal{A}),$ (9)
with $c>0$. Then, for all $x\in\Omega$
$\frac{\mathrm{d}P}{\mathrm{d}Q}(x)$ =$~{}\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{Q}\frac{1}{c},\mbox{ and }\frac{\mathrm{d}Q}{\mathrm{d}P}(x)$ =$\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{P}c.$ (10)
Proof:
First, note that (9) implies that the measures $P$ and $Q$ are mutually absolutely continuous. Hence, from Theorem 1, it follows that for all $\mathcal{A}\in\mathscr{F}$,
$P\left(\mathcal{A}\right)=\int_{\mathcal{A}}\mathrm{d}P(x)=\int_{\mathcal{A}}\frac{\mathrm{d}P}{\mathrm{d}Q}(x)\;\mathrm{d}Q(x).$ (11)
On the other hand, the equality (9) also implies
$P\left(\mathcal{A}\right)=\frac{1}{c}Q(\mathcal{A})=\int_{\mathcal{A}}\frac{1}{c}\mathrm{d}Q(x).$ (12)
Hence, it follows directly from Theorem 1 that the Radon-Nikodym derivative $\frac{\mathrm{d}P}{\mathrm{d}Q}$ is unique almost surely with respect to $Q$. Thus, $\frac{\mathrm{d}P}{\mathrm{d}Q}(x)\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{Q}$$\frac{1}{c}$. Using similar arguments and the fact that $P$ and $Q$ are mutually absolutely continuous, it is verified that $\frac{\mathrm{d}Q}{\mathrm{d}P}(x)\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{P}$$c$. ∎
In the case in which $c=1$ in (10), measures $P$ and $Q$ are identical, thus, $\frac{\mathrm{d}P}{\mathrm{d}Q}(x)\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{Q}$ $\frac{\mathrm{d}Q}{\mathrm{d}P}(x)\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{P}$ $1$
The following folklore theorem is often referred to as the “chain rule”.
Theorem 4 (Chain Rule)
Let $P$, $Q$, and $R$ be three measures on the measurable space $\left(\Omega,\mathscr{F}\right)$ such that ${P}\ll{Q}$; ${Q}\ll{R}$; and $Q$ and $R$ are $\sigma$-finite measures. Then,
$\frac{\mathrm{d}P}{\mathrm{d}R}(x)\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{R}\frac{\mathrm{d}P}{\mathrm{d}Q}(x)\frac{\mathrm{d}Q}{\mathrm{d}R}(x).$ (13)
Proof:
From the assumptions of the theorem, it follows that for all $\mathcal{A}\in\mathscr{F}$,
$P(\mathcal{A})$ $=$ $\int_{\mathcal{A}}\mathrm{d}P(x)=\int_{\mathcal{A}}\frac{\mathrm{d}P}{\mathrm{d}Q}(x)\mathrm{d}Q(x)$ (14)
$=$ $\int_{\mathcal{A}}\frac{\mathrm{d}P}{\mathrm{d}Q}(x)\frac{\mathrm{d}Q}{\mathrm{d}R}(x)\mathrm{d}R(x)$ (15)
$=$ $\int_{\mathcal{A}}\frac{\mathrm{d}P}{\mathrm{d}R}(x)\mathrm{d}R(x),$ (16)
where the second equality in (14) follows from Theorem 1; and the equality in (15) follows from Theorem 2 . The equality in (16) holds from Theorem 1 and by noticing that $P~{}\ll~{}R$. Therefore, the equalities in (15) and (16) together with Theorem 1 imply (13), which completes the proof. ∎
The following folklore theorem shows the connection between the Radon-Nikodym derivative and its multiplicative inverse.
Theorem 5 (Multiplicative Inverse)
Let $P$ and $Q$ be two mutually absolutely continuous measures on the measurable space $\left(\Omega,\mathscr{F}\right)$; and assume that for all $x\in\Omega$, $\frac{\mathrm{d}Q}{\mathrm{d}P}(x)>0$. Then,
$\frac{\mathrm{d}P}{\mathrm{d}Q}(x)$ =$\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{Q}$ $\left(\frac{\mathrm{d}Q}{\mathrm{d}P}(x)\right)^{-1}.$ (17)
Proof:
From Theorem 4, it follows that
$\frac{\mathrm{d}P}{\mathrm{d}Q}(x)\frac{\mathrm{d}Q}{\mathrm{d}P}(x)\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{Q}\frac{\mathrm{d}Q}{\mathrm{d}Q}(x)\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{Q}1,$ (18)
where the last equality follows from Theorem 3, with $c~{}=~{}1$. This completes the proof. ∎
The subsequent folklore theorem establishes the linearity of the Radon-Nikodym derivative.
Theorem 6 (Linearity)
Let $P$ be a $\sigma$-finite measure on $\left(\Omega,\mathscr{F}\right)$ and let also $Q_{1},Q_{2},\ldots,Q_{n}$ be finite measures on $\left(\Omega,\mathscr{F}\right)$ absolutely continuous with respect to $P$. Let $c_{1},c_{2},\ldots,c_{n}$ be positive reals; and let $S$ be a finite measure on $\left(\Omega,\mathscr{F}\right)$ such that for all $\mathcal{A}\in\mathscr{F}$, $S(\mathcal{A})=\sum_{t=1}^{n}c_{t}Q_{t}(\mathcal{A})$. Then,
$\frac{\mathrm{d}S}{\mathrm{d}P}(x)$ =$\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{P}$ $\sum_{t=1}^{n}c_{t}\frac{\mathrm{d}Q_{t}}{\mathrm{d}P}(x).$ (19)
Proof:
The proof starts by noticing that, from the assumptions of the theorem, it holds that ${S}\ll{P}$. Hence, for all $\mathcal{A}\in\mathscr{F}$, it holds that
$\int_{\mathcal{A}}\frac{\mathrm{d}S}{\mathrm{d}P}(x)\mathrm{d}P(x)=\int_{\mathcal{A}}\mathrm{d}S(x)=\sum_{t=1}^{n}c_{t}Q_{t}(\mathcal{A})$ (20)
$=$ $\sum_{t=1}^{n}\int_{\mathcal{A}}c_{t}\mathrm{d}Q_{t}(x)=\sum_{t=1}^{n}\int_{\mathcal{A}}c_{t}\frac{\mathrm{d}Q_{t}}{\mathrm{d}P}(x)\mathrm{d}P(x)$ (21)
$=$ $\int_{\mathcal{A}}\sum_{t=1}^{n}c_{t}\frac{\mathrm{d}Q_{t}}{\mathrm{d}P}(x)\mathrm{d}P(x),$ (22)
where the first equality in (20) and the last equality in (21) follow from Theorem 2; and the equality (22) follows from the additivity property of the integral [5, Corollary 1.6.4]. The proof ends by using Theorem 1, which implies the equality in (19) from (22). ∎
The following folklore theorem establishes the continuity of the Radon-Nikodym derivative.
Theorem 7 (Continuity)
Let $P$ be a $\sigma$-finite measure on $\left(\Omega,\mathscr{F}\right)$, and let $Q_{1},Q_{2},\cdots$ be an infinite sequence of $\sigma$-finite measures on $\left(\Omega,\mathscr{F}\right)$, converging to a measure $Q$. Suppose that for all $n\in\mathbb{N}$, ${Q_{n}}\ll{P}$. Then, ${Q}\ll{P}$ and
$\lim_{n\rightarrow\infty}\frac{\mathrm{d}Q_{n}}{\mathrm{d}P}(x)\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{P}\frac{\mathrm{d}Q}{\mathrm{d}P}(x).$ (23)
Proof:
From the assumptions of the theorem, for all $\mathcal{A}\in\mathscr{F}$, it holds that
$Q(\mathcal{A})$ $=$ $\lim_{n\rightarrow\infty}Q_{n}(\mathcal{A})$ (24)
$=$ $\lim_{n\rightarrow\infty}\int_{\mathcal{A}}\frac{\mathrm{d}Q_{n}}{\mathrm{d}P}(x)\mathrm{d}P(x)$ (25)
$=$ $\int_{\mathcal{A}}\lim_{n\to\infty}\frac{\mathrm{d}Q_{n}}{\mathrm{d}P}(x)\mathrm{d}P(x),$ (26)
where the equality in (25) follows from Theorem 2 ; and the equality in (26) follows from [5, Theorem 1.6.2]. The equality in (26) implies that ${Q}\ll{P}$. Hence, for all $\mathcal{A}\in\mathscr{F}$, it holds that
$Q(\mathcal{A})=\int_{\mathcal{A}}\frac{\mathrm{d}Q}{\mathrm{d}P}(x)\mathrm{d}P(x).$ (27)
Therefore, the equalities in (26) and (27) jointly with Theorem 1 imply equation (23), which completes the proof. ∎
The ensuing folklore theorem establishes the relation between the Radon-Nikodym derivative of a product measure with respect to its component measures.
Theorem 8 (Product of Measures)
For all $i\in\{1,2\}$, let $P_{i}$ and $Q_{i}$ be a finite and a $\sigma$-finite measure on $\left(\Omega_{i},\mathscr{F}_{i}\right)$, respectively; with ${P_{i}}\ll{Q_{i}}$. Let also $P_{1}P_{2}$ and $Q_{1}Q_{2}$ be the product measures on $\left(\Omega_{1}\times\Omega_{2},\mathscr{F}_{1}\times\mathscr{F}_{2}\right)$ formed by $P_{1}$ and $P_{2}$; and $Q_{1}$ and $Q_{2}$, respectively. Then,
$\frac{\mathrm{d}P_{1}P_{2}}{\mathrm{d}Q_{1}Q_{2}}\left(x_{1},x_{2}\right)$ =$\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{Q_{1}Q_{2}}$ $\frac{\mathrm{d}P_{1}}{\mathrm{d}Q_{1}}\left(x_{1}\right)\frac{\mathrm{d}P_{2}}{\mathrm{d}Q_{2}}\left(x_{2}\right).$ (28)
Proof:
From the assumptions of the theorem, for all $\mathcal{A}\in\left(\Omega_{1}\times\Omega_{2}\right)$,
$P_{1}P_{2}\left(\mathcal{A}\right)$ $=$ $\int_{\mathcal{A}}\mathrm{d}P_{1}P_{2}\left(x_{1},x_{2}\right)$ (29)
$=$ $\int\int_{\mathcal{A}_{x_{2}}}\mathrm{d}P_{1}\left(x_{1}\right)\mathrm{d}P_{2}\left(x_{2}\right)$ (30)
$=$ $\int\int_{\mathcal{A}_{x_{2}}}\dfrac{\mathrm{d}P_{1}\left(x_{1}\right)}{\mathrm{d}Q_{1}}\mathrm{d}Q_{1}\left(x_{1}\right)\mathrm{d}P_{2}\left(x_{2}\right)$ (31)
$=$ $\int\int_{\mathcal{A}_{x_{2}}}\dfrac{\mathrm{d}P_{1}\left(x_{1}\right)}{\mathrm{d}Q_{1}}\dfrac{\mathrm{d}P_{2}\left(x_{2}\right)}{\mathrm{d}Q_{2}}\mathrm{d}Q_{1}\left(x_{1}\right)\mathrm{d}Q_{2}\left(x_{2}\right)$ (32)
$=$ $\int_{\mathcal{A}}\dfrac{\mathrm{d}P_{1}}{\mathrm{d}Q_{1}}\left(x_{1}\right)\dfrac{\mathrm{d}P_{2}}{\mathrm{d}Q_{2}}\left(x_{2}\right)\mathrm{d}Q_{1}Q_{2}\left(x_{1},x_{2}\right),$ (33)
where $\mathcal{A}_{x_{2}}$ is the section of the set $\mathcal{A}$ determined by $x_{2}$, namely, $\mathcal{A}_{x_{2}}\triangleq\left\{x_{1}\in\Omega_{1}:(x_{1},x_{2})\in\mathcal{A}\right\}$; the equality in (29) arises from the definition of $P_{1}P_{2}$ as the product of $P_{1}$ and $P_{2}$; the equality in (31) is a direct consequence of Theorem 2; the equality in (32) follows from Theorem 1; and finally, the equality in (33) is due to the construction of $Q_{1}Q_{2}$ as the product measure of $Q_{1}$ and $Q_{2}$.
The proof follows by observing that from the equality in (33), it holds that ${P_{1}P_{2}}\ll{Q_{1}Q_{2}}$. Thus, for all $\mathcal{A}~{}\in~{}\mathscr{F}_{1}~{}\times~{}\mathscr{F}_{2}$,
$P_{1}P_{2}\left(\mathcal{A}\right)$ $=$ $\int_{\mathcal{A}}\dfrac{\mathrm{d}P_{1}P_{2}}{\mathrm{d}Q_{1}Q_{2}}\left(x_{1},x_{2}\right)\mathrm{d}Q_{1}Q_{2}\left(x_{1},x_{2}\right).$ (34)
The equalities in (33) and (34), together with Theorem 1, imply the equality in (28), which completes the proof. ∎
IV Advanced folklore theorems
This section requires some additional notation. In particular, denote by $\triangle\left(\mathcal{X},\mathscr{F}_{\mathcal{X}}\right)$, or simply $\triangle\left(\mathcal{X}\right)$, the set of all probability measures on the measurable space $\left(\mathcal{X},\mathscr{F}_{\mathcal{X}}\right)$, where $\mathscr{F}_{\mathcal{X}}$ is a $\sigma$-algebra on $\mathcal{X}$. Using this notation, conditional probability measures can be defined as follows.
Definition 1 (Conditional Probability)
A family $P_{Y|X}~{}\triangleq~{}(P_{Y|X=x})_{x\in\mathcal{X}}$ of elements of $\triangle(\mathcal{Y},\mathscr{F}_{\mathcal{Y}})$ indexed by $\mathcal{X}$ is said to be a conditional probability measure if, for all sets $\mathcal{A}\in\mathscr{F}_{\mathcal{Y}}$, the map
$\mathcal{X}$$\to$ $[0,1]$ (35)
$x$ $\mapsto$ $P_{Y|X=x}(\mathcal{A})$ (36)
is Borel measurable. The set of such conditional probability measures is denoted by $\triangle(\mathcal{Y}|\mathcal{X})$.
A conditional probability $P_{Y|X}\in\triangle(\mathcal{Y}|\mathcal{X})$ and a probability measure $P_{X}\in\triangle(\mathcal{X})$ determine two unique probability measures in $\triangle(\mathcal{X}\times\mathcal{Y})$ and $\triangle(\mathcal{Y}\times\mathcal{X})$, respectively. These probability measures are denoted by $P_{XY}$ and $P_{YX}$, respectively, and for all sets $\mathcal{A}\in\mathscr{F}_{\mathcal{X}}\times\mathscr{F}_{\mathcal{Y}}$, it follows that
$P_{XY}\left(\mathcal{A}\right)$ $=$ $\int P_{Y|X=x}\left(\mathcal{A}_{x}\right)\mathrm{d}P_{X}\left(x\right),$ (37)
where $\mathcal{A}_{x}$ is the section of the set $\mathcal{A}$ determined by $x$, namely,
$\mathcal{A}_{x}\triangleq\left\{y\in\mathcal{Y}:(x,y)\in\mathcal{A}\right\}.$ (38)
Alternatively, for all sets $\mathcal{B}\in\mathscr{F}_{\mathcal{Y}}\times\mathscr{F}_{\mathcal{X}}$, it follows that
$P_{YX}\left(\mathcal{B}\right)$ $=$ $\int P_{Y|X=x}\left(\mathcal{B}_{x}\right)\mathrm{d}P_{X}\left(x\right),$ (39)
where $\mathcal{B}_{x}$ is the section of the set $\mathcal{B}$ determined by $x$. For all sets $\mathcal{A}\in\mathscr{F}_{\mathcal{X}}\times\mathscr{F}_{\mathcal{Y}}$, let the set $\hat{\mathcal{A}}\in\mathscr{F}_{\mathcal{Y}}\times\mathscr{F}_{\mathcal{X}}$ be such that
$\hat{\mathcal{A}}$ $=$ $\left\{(y,x)\in\mathcal{Y}\times\mathcal{X}:(x,y)\in\mathcal{A}\right\}.$ (40)
Then, from (37) and (39), it holds that
$P_{XY}\left(\mathcal{A}\right)$ $=$ $P_{YX}\left(\hat{\mathcal{A}}\right).$ (41)
Using this notation, the notion of marginal probability measures can be introduced as follows.
Definition 2 (Marginals)
Given two joint probability measures $P_{XY}\in\triangle\left(\mathcal{X}\times\mathcal{Y}\right)$ and $P_{YX}\in\triangle\left(\mathcal{Y}\times\mathcal{X}\right)$, satisfying (41), the marginal probability measures in $\triangle\left(\mathcal{X}\right)$ and $\triangle\left(\mathcal{Y}\right)$, denoted by $P_{X}$ and $P_{Y}$, respectively satisfy for all sets $\mathcal{A}\in\mathscr{F}_{\mathcal{X}}$ and for all sets $\mathcal{B}\in\mathscr{F}_{\mathcal{Y}}$,
$P_{X}\left(\mathcal{A}\right)$ $\triangleq P_{XY}\left(\mathcal{A}\times\mathcal{Y}\right)=$ $P_{YX}\left(\mathcal{Y}\times\mathcal{A}\right);\mbox{ and }$ (42)
$P_{Y}\left(\mathcal{B}\right)$ $\triangleq P_{XY}\left(\mathcal{X}\times\mathcal{B}\right)=$ $P_{YX}\left(\mathcal{B}\times\mathcal{X}\right).$ (43)
From the total probability theorem [5, Theorem 4.5.2], it follows that for all $\mathcal{A}\in\mathscr{F}_{\mathcal{Y}}$,
$P_{Y}(\mathcal{A})=\int\int_{\mathcal{A}}\mathrm{d}P_{Y|X=x}(y)\mathrm{d}P_{X}(x);$ (44)
and for all $\mathcal{B}\in\mathscr{F}_{\mathcal{X}}$,
$P_{X}(\mathcal{B})=\int\int_{\mathcal{B}}\mathrm{d}P_{X|Y=y}(x)\mathrm{d}P_{Y}(y).$ (45)
The joint probability measures $P_{XY}$ and $P_{YX}$ can be described via the conditional probability measure $P_{Y|X}$ and the probability measure $P_{X}$ as in (37) and in (39); or via the conditional probability measure $P_{X|Y}\in\triangle\left(\mathcal{X}|\mathcal{Y}\right)$ and the marginal probability measure $P_{Y}\in\triangle\left(\mathcal{Y}\right)$. More specifically, for all sets $\mathcal{A}\in\mathscr{F}_{\mathcal{X}}\times\mathscr{F}_{\mathcal{Y}}$, it follows that
$P_{XY}\left(\mathcal{A}\right)$ $=$ $\int P_{X|Y=y}\left(\mathcal{A}_{y}\right)\mathrm{d}P_{Y}\left(y\right),$ (46)
where $\mathcal{A}_{y}$ is the section of the set $\mathcal{A}$ determined by $y$, namely,
$\mathcal{A}_{y}\triangleq\left\{x\in\mathcal{X}:(x,y)\in\mathcal{A}\right\}.$ (47)
Alternatively, for all sets $\mathcal{B}\in\mathscr{F}_{\mathcal{Y}}\times\mathscr{F}_{\mathcal{X}}$, it follows that
$P_{YX}\left(\mathcal{B}\right)$ $=$ $\int P_{X|Y=y}\left(\mathcal{B}_{y}\right)\mathrm{d}P_{Y}\left(y\right),$ (48)
where $\mathcal{B}_{y}$ is the section set of $\mathcal{B}$ determined by $y$.
Within this context, the following folklore theorem highlights a property of conditional measures, which is reminiscent of the unit measure axiom in probability theory.
Theorem 9 (Unit Measure)
Consider the conditional probability measures $P_{Y|X}\in\triangle\left(\mathcal{Y}|\mathcal{X}\right)$ and $P_{X|Y}\in\triangle\left(\mathcal{X}|\mathcal{Y}\right)$; the probability measures $P_{Y}\in\triangle\left(\mathcal{Y}\right)$ and $P_{X}\in\triangle\left(\mathcal{X}\right)$ that satisfy (44) and (45). Assume that for all $x\in\mathcal{X}$, the probability measure ${P_{Y|X=x}}\ll{P_{Y}}$. Then,
$\int\frac{\mathrm{d}P_{Y|X=x}}{\mathrm{d}P_{Y}}(y)\mathrm{d}P_{X}(x)$ =$\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{P_{Y}}$ $1$ (49)
Proof:
For all $\mathcal{A}\in\mathscr{F}_{\mathcal{Y}}$, from (44), it holds that
$P_{Y}(\mathcal{A})$ $=$ $\int\int_{\mathcal{A}}\mathrm{d}P_{Y|X=x}(y)\mathrm{d}P_{X}(x)$ (50)
$=$ $\int\int_{\mathcal{A}}\frac{\mathrm{d}P_{Y|X=x}}{\mathrm{d}P_{Y}}(y)\mathrm{d}P_{Y}(y)\mathrm{d}P_{X}(x)$ (51)
$=$ $\int_{\mathcal{A}}\int\frac{\mathrm{d}P_{Y|X=x}}{\mathrm{d}P_{Y}}(y)\mathrm{d}P_{X}(x)\mathrm{d}P_{Y}(y),$ (52)
where the equality in (51) follows from a change of measure (Theorem 2). Moreover, (52) is obtained using Fubini’s theorem [5, Theorem 2.6.6]. The proof proceeds by noticing that $P_{Y}(\mathcal{A})=\int_{\mathcal{A}}\mathrm{d}P_{Y}(y)$, and thus from Theorem 1 and the equality in (52), the statement in (49) holds. ∎
The following folklore theorem is reminiscent of the Bayes rule.
Theorem 10 (Bayes-like rule)
Consider the conditional probability measures $P_{Y|X}$ and $P_{X|Y}$; the probability measures $P_{Y}$ and $P_{X}$ that satisfy (44) and (45); and the joint probability measures $P_{YX}$ and $P_{XY}$ in (39) and (46) respectively. Let also $P_{X}P_{Y}\in\triangle\left(\mathcal{X}\times\mathcal{Y},\mathscr{F}_{\mathcal{X}}\times\mathscr{F}_{\mathcal{Y}}\right)$ and $P_{Y}P_{X}\in\triangle\left(\mathcal{Y}\times\mathcal{X},\mathscr{F}_{\mathcal{Y}}\times\mathscr{F}_{\mathcal{X}}\right)$ be the measures formed by the product of the marginals $P_{X}$ and $P_{Y}$. Assume that:
  • $\left(a\right)$
    For all $x\in\mathcal{X}$${P_{Y|X=x}}\ll{P_{Y}}$; and
  • $\left(b\right)$
    For all $y\in\mathcal{Y}$${P_{X|Y=y}}\ll{P_{X}}$.
Then,
$\frac{\mathrm{d}P_{XY}}{\mathrm{d}P_{X}P_{Y}}\left(x,y\right)$ =$\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{P_{X}P_{Y}}$ $\frac{\mathrm{d}P_{X|Y=y}}{\mathrm{d}P_{X}}\left(x\right)$ (53)
=$\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{P_{X}P_{Y}}$ $\frac{\mathrm{d}P_{Y|X=x}}{\mathrm{d}P_{Y}}\left(y\right)$ (54)
=$\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{P_{X}P_{Y}}$ $\frac{\mathrm{d}P_{YX}}{\mathrm{d}P_{Y}P_{X}}\left(y,x\right).$ (55)
Proof:
Note that assumptions $\left(a\right)$ and $\left(b\right)$ are sufficient for the Radon-Nikodym derivatives of $P_{XY}$ with respect to $P_{X}P_{Y}$ and $P_{YX}$ with respect to $P_{Y}P_{X}$ to exist. Hence, it follows that for all sets $\mathcal{A}~{}\in~{}\mathscr{F}_{\mathcal{X}}\times\mathscr{F}_{\mathcal{Y}}$,
$P_{XY}\left(\mathcal{A}\right)$ $=$ $\int_{\mathcal{A}}\frac{\mathrm{d}P_{XY}}{\mathrm{d}P_{X}P_{Y}}(x,y)\mathrm{d}P_{X}P_{Y}(x,y),$ (56)
which follows from Theorem 2. Note also that from (46), it follows that
$P_{XY}\left(\mathcal{A}\right)$ $=$ $\int\int_{\mathcal{A}_{y}}\mathrm{d}P_{X|Y=y}\left(x\right)\mathrm{d}P_{Y}\left(y\right)$ (57)
$=$ $\int\int_{\mathcal{A}_{y}}\frac{\mathrm{d}P_{X|Y=y}}{\mathrm{d}P_{X}}\left(x\right)\mathrm{d}P_{X}\left(x\right)\mathrm{d}P_{Y}\left(y\right)$ (58)
$=$ $\int\int\mathds{1}_{\mathcal{A}_{y}}(x)\frac{\mathrm{d}P_{X|Y=y}}{\mathrm{d}P_{X}}\left(x\right)\mathrm{d}P_{X}\left(x\right)\mathrm{d}P_{Y}\left(y\right)$ (59)
$=$ $\int\mathds{1}_{\mathcal{A}}(x,y)\frac{\mathrm{d}P_{X|Y=y}}{\mathrm{d}P_{X}}\left(x\right)\mathrm{d}P_{X}P_{Y}\left(x,y\right)$ (60)
$=$ $\int_{\mathcal{A}}\frac{\mathrm{d}P_{X|Y=y}}{\mathrm{d}P_{X}}\left(x\right)\mathrm{d}P_{X}P_{Y}\left(x,y\right),$ (61)
where, the set $\mathcal{A}_{y}$ is defined in (47). Moreover, the equality in (58) follows from Assumption $\left(b\right)$ and Theorem 1. Similarly, from (37), it follows that
$P_{XY}\left(\mathcal{A}\right)$ $=$ $\int\int_{\mathcal{A}_{x}}\mathrm{d}P_{Y|X=x}\left(y\right)\mathrm{d}P_{X}\left(x\right)$ (62)
$=$ $\int\int_{\mathcal{A}_{x}}\frac{\mathrm{d}P_{Y|X=x}}{\mathrm{d}P_{Y}}\left(y\right)\mathrm{d}P_{Y}\left(y\right)\mathrm{d}P_{X}\left(x\right)$ (63)
$=$ $\int\int\mathds{1}_{\mathcal{A}_{x}}(y)\frac{\mathrm{d}P_{Y|X=x}}{\mathrm{d}P_{Y}}\left(y\right)\mathrm{d}P_{Y}\left(y\right)\mathrm{d}P_{X}\left(x\right)$ (64)
$=$ $\int\int\mathds{1}_{\mathcal{A}_{y}}(x)\frac{\mathrm{d}P_{Y|X=x}}{\mathrm{d}P_{Y}}\left(y\right)\mathrm{d}P_{X}\left(x\right)\mathrm{d}P_{Y}\left(y\right)$ (65)
$=$ $\int\mathds{1}_{\mathcal{A}}(x,y)\frac{\mathrm{d}P_{Y|X=x}}{\mathrm{d}P_{Y}}\left(y\right)\mathrm{d}P_{X}P_{Y}\left(x,y\right)$ (66)
$=$ $\int_{\mathcal{A}}\frac{\mathrm{d}P_{Y|X=x}}{\mathrm{d}P_{Y}}\left(y\right)\mathrm{d}P_{X}P_{Y}\left(x,y\right),$ (67)
where, the set $\mathcal{A}_{x}$ is defined in (38). Moreover, the equality in (63) follows from Assumption $\left(a\right)$ and Theorem 1; and the equality in (65) follows by exchanging the order of integration [5, Theorem 2.6.6]. Finally, from (41), it follows that
$P_{XY}\left(\mathcal{A}\right)=\int_{\hat{\mathcal{A}}}\mathrm{d}P_{YX}\left(y,x\right)$ (68)
$=$ $\int_{\hat{\mathcal{A}}}\frac{\mathrm{d}P_{YX}}{\mathrm{d}P_{Y}P_{X}}\left(y,x\right)\mathrm{d}P_{Y}P_{X}\left(y,x\right)$
$=$ $\int\mathds{1}_{\hat{\mathcal{A}}}(y,x)\frac{\mathrm{d}P_{YX}}{\mathrm{d}P_{Y}P_{X}}\left(y,x\right)\mathrm{d}P_{Y}P_{X}\left(y,x\right)$ (69)
$=$ $\int\int\mathds{1}_{\hat{\mathcal{A}}_{x}}(y)\frac{\mathrm{d}P_{YX}}{\mathrm{d}P_{Y}P_{X}}\left(y,x\right)\mathrm{d}P_{Y}(y)\mathrm{d}P_{X}(x)$ (70)
$=$ $\int\int\mathds{1}_{\hat{\mathcal{A}}_{y}}(x)\frac{\mathrm{d}P_{YX}}{\mathrm{d}P_{Y}P_{X}}\left(y,x\right)\mathrm{d}P_{X}(x)\mathrm{d}P_{Y}(y)$ (71)
$=$ $\int\mathds{1}_{\hat{\mathcal{A}}}(x,y)\frac{\mathrm{d}P_{YX}}{\mathrm{d}P_{Y}P_{X}}\left(y,x\right)\mathrm{d}P_{X}P_{Y}(x,y)$ (72)
$=$ $\int_{\mathcal{A}}\frac{\mathrm{d}P_{YX}}{\mathrm{d}P_{Y}P_{X}}\left(y,x\right)\mathrm{d}P_{X}P_{Y}(x,y),$ (73)
where the set $\hat{\mathcal{A}}$ is defined in (40). By performing a change of measure using Theorem 2 and assumption $\left(a\right)$, the equality in (68) is obtained; and the equality in (71) follows by exchanging the order of integration [5, Theorem 2.6.6].
The proof is completed from Theorem 1 and by combining equations (56),  (61),  (67) and (73), which establish (53), (54) and (55). ∎
Theorem 11 (Inverse Bayes-like Rule)
Consider the conditional probability measures $P_{Y|X}$ and $P_{X|Y}$; and the probability measures $P_{Y}$ and $P_{X}$ that satisfy (44) and (45); and the joint probability measures $P_{YX}$ and $P_{XY}$ in (39) and (46) respectively. Assume that:
  • $\left(a\right)$
    For all $x\in\mathcal{X}$${P_{Y}}\ll{P_{Y|X=x}}$; and
  • $\left(b\right)$
    For all $y\in\mathcal{Y}$${P_{X}}\ll{P_{X|Y=y}}$.
Then,
$\frac{\mathrm{d}P_{X}P_{Y}}{\mathrm{d}P_{XY}}\left(x,y\right)\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{P_{XY}}\frac{\mathrm{d}P_{X}}{\mathrm{d}P_{X|Y=y}}\left(x\right)$ (74)
=$\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{P_{XY}}$ $\frac{\mathrm{d}P_{Y}}{\mathrm{d}P_{Y|X=x}}\left(y\right)\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{P_{YX}}\frac{\mathrm{d}P_{Y}P_{X}}{\mathrm{d}P_{YX}}\left(y,x\right).$
Proof:
The proof follows along the same lines as the proof of Theorem 10. ∎
V Conclusions and Final Remarks
This paper provides proofs for several well-known folklore theorems, as well as some other lesser-known results in information theory involving the Radon-Nikodym derivative. Notably, these theorems serve as fundamental tools for establishing various properties of Shannon’s information measures in a unified framework, that is, regardless of whether the underlying measures admit a probability mass function or a probability density function. Nonetheless, the proof of these properties are left out of the scope of this paper, due to space constraints. Finally, it is important to highlight that the presented results hold in full generality, as the reference measure is assumed to be any $\sigma$-finite measure.

### v2

@@
 Abstract
-Rigorous statements and formal proofs are presented for both foundational and advanced folklore theorems on the Radon-Nikodym derivative.
-The cases of product and marginal measures are carefully considered; and the hypothesis under which the statements hold are rigorously enumerated.
+In this paper, rigorous statements and formal proofs are presented for both foundational and advanced folklore theorems on the Radon-Nikodym derivative.
+The cases of conditional and marginal probability measures are carefully considered, which leads to an identity involving the sum of mutual and lautum information suggesting a new interpretation for such a sum.
 I Introduction
@@
 The existence of such folklore theorems in this area arises in part from the fact that all Shannon’s information measures can be defined in terms of the RND.
-Interestingly, most properties of the RND are often presented as comments in most textbooks on measure theory and probability theory, c.f, [5, 6, 7, 8, 9, 10, 11]
+Interestingly, most properties of the RND are often presented without proof in most textbooks on measure theory and probability theory, c.f, [5, 6, 7, 8, 9, 10, 11], with the exception of some properties presented in [12].
 Claude Shannon did not use the RND in his foundational papers [13, 14] to define entropy and mutual information.
-Instead, Shannon opted for restricting his publications to the case in which measures either possess a probability mass function or a probability density function, which are both instances of RNDs.
+Instead, Shannon opted for restricting his work to the case in which measures either possess a probability mass function or a probability density function, which are both instances of RNDs.
 This choice significantly influenced the presentation of most subsequent results in information theory and established the style in which classical textbooks were written [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31].
@@
 That is, independently of the measure used as a reference, e.g., the counting measure, the Lesbegue measure, etc.
-Some recent results whose presentation relies on the RND are for instance [33, 34, 35, 36, 37, 38, 39, 40, 41, 42].
-In the light of the central role of the RND in information theory, this paper presents rigorous statements and formal proofs for the most common folklore theorems on the RND.
+Some recent results whose presentation relies on the RND are for instance [34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45].
+At the light of this observation, this paper presents rigorous statements and alternative proofs for the most common folklore theorems on the RND.
 These include the change of measure theorem, the proportional-measures theorem, the chain rule, the multiplicative inverse, linearity and continuity theorems, as well as the product-measure theorem.
-Less common theorems such as the unit measure and two Bayes-like rules of the RND are also formally proved.
-This is the first time such material appears in peer-reviewed literature.
-Finally, it is important to highlight that particular attention has been put on rigorously stating the conditions under which these theorems hold, hopefully providing a valuable reference for researchers and students in information theory.
+Less common theorems such as the unit measure and two Bayes-like rules of the RND are also formally proved for the first time in peer-reviewed literature.
+Finally, these novel theorems are used to prove a key identity involving the sum of mutual and lautum information, which is often observed in statistical learning [46, 42].
+Hopefully, these contributions would provide a valuable reference for researchers and students in information theory.
 II Preliminaries
@@
 Using this notation, the Radon-Nikodym derivative is introduced by the following theorem.
-Theorem 1 (Radon-Nikodym theorem, [5, Theorem 2.2.1])
-Let $P$ and $Q$ be two measures on a given measurable space $\left(\Omega,\mathscr{F}\right)$, such that $Q$ is $\sigma$-finite and ${P}\ll{Q}$.
-Then, there exists a nonnegative Borel measurable function $g:~{}\Omega~{}\to~{}\mathbb{R}$ such that for all $\mathcal{A}\in\mathscr{F}$,
+Theorem 1 (Radon-Nikodym theorem, [5, Theorem 2.2.1]).
+Let $P$ be a signed measure and $Q$ be a $\sigma$-finite measure both on a given measurable space $\left(\Omega,\mathscr{F}\right)$, with ${P}\ll{Q}$.
+Then, there exists a Borel measurable function $g:\leavevmode\nobreak\ \Omega\leavevmode\nobreak\ \to\leavevmode\nobreak\ \mathbb{R}$ such that for all $\mathcal{A}\in\mathscr{F}$,
 $P(\mathcal{A})=\int_{\mathcal{A}}g(x)\mathrm{d}Q(x).$ (1)
 Moreover, if another function $h$ satisfies for all $\mathcal{A}\in\mathscr{F}$ that $P(\mathcal{A})=\int_{\mathcal{A}}h(x)\mathrm{d}Q(x)$, then $g(x)\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{Q}$$h(x)$.
-The function $g$ in (1) is often referred to as the Radom-Nikodym derivative of $P$ with respect to $Q$; and is also written as $\frac{\mathrm{d}P}{\mathrm{d}Q}$, such that $g(x)~{}=~{}\frac{\mathrm{d}P}{\mathrm{d}Q}(x).$
-The Radon-Nikodym theorem is the foundational tool from which many folklore theorems in information theory originate. Some of these folklore theorems are thoroughly studied in the following sections.
+The function $g$ in (1) is often referred to as the Radom-Nikodym derivative of $P$ with respect to $Q$; and is also written as $\frac{\mathrm{d}P}{\mathrm{d}Q}$, such that $g(x)\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{\mathrm{d}P}{\mathrm{d}Q}(x).$
+The Radon-Nikodym derivative exhibits a number of properties from which many foundational results originate.
+Some of these properties are certainly folklore theorems and thus, the purpose of the following section is to provide formal proofs.
+Other properties are formally proved in literature and thus, such properties are simply stated for the sake of completeness.
+In some cases, alternative proofs are presented, for privileging simplicity.
 III Basic Folklore Theorems
 This section focuses on basic folklore theorems, where “basic” denotes their well-established nature. One of the most common folklore theorems is often referred to as the “change of measure” theorem.
-Theorem 2 (Change of Measure)
-Let $P$ and $Q$ be two measures on the measurable space $\left(\Omega,\mathscr{F}\right)$ with ${P}\ll{Q}$; and $Q$ a $\sigma$-finite measure.
-Let $f:\Omega\to\mathds{R}$ be a Borel measurable function such that the integral $\int_{\Omega}f(x)\mathrm{d}P(x)$ exists.
+Theorem 2 (Change of Measure).
+Let $P$ be a signed measure and $Q$ be a $\sigma$-finite measure both on a given measurable space $\left(\Omega,\mathscr{F}\right)$, with ${P}\ll{Q}$.
+Let $f:\Omega\to\mathbb{R}$ be a Borel measurable function such that the integral $\int_{\Omega}f(x)\mathrm{d}P(x)$ exists.
 Then, for all $\mathcal{A}\in\mathscr{F}$,
 $\int_{\mathcal{A}}f(x)\mathrm{d}P(x)=\displaystyle\int_{\mathcal{A}}f(x)\frac{\mathrm{d}P}{\mathrm{d}Q}(x)\mathrm{d}Q(x).$ (2)
+A proof of this theorem is presented in [12, Proposition 3.9]. For the sake of completeness, an alternative proof is presented hereunder, using the previously described notation.
 Proof:
-[…]
+Assume that $f$ is a simple function, i.e.
+for all $x\in\Omega$, $f(x)=\sum_{i=1}^{m}a_{i}\mathds{1}_{\mathcal{A}_{i}}(x),$ for some finite $m\in\mathbb{N}$, disjoint sets $\mathcal{A}_{1},\ldots,\mathcal{A}_{m}$ in $\mathscr{F}$, and reals $a_{1},\ldots,a_{m}$.
+For all $\mathcal{A}\in\mathscr{F}$ and $i\in\{1,\ldots,m\}$, let $\mathcal{B}_{i}=\mathcal{A}\cap\mathcal{A}_{i}$.
+Then,
 $\int_{\mathcal{A}}f(x)\frac{\mathrm{d}P}{\mathrm{d}Q}(x)\mathrm{d}Q(x)=\int_{\mathcal{A}}\frac{\mathrm{d}P}{\mathrm{d}Q}(x)\sum_{i=1}^{n}a_{i}\mathds{1}_{\mathcal{A}_{i}}(x)\mathrm{d}Q(x)$ (3)
-$=$ $\sum_{i=1}^{n}a_{i}\int_{\mathcal{B}_{i}}\frac{\mathrm{d}P}{\mathrm{d}Q}(x)\mathrm{d}Q(x)$ (4)
-$=$ $\sum_{i=1}^{n}a_{i}P(\mathcal{B}_{i}),$ (5)
-where the equality in (LABEL:EqMartinezPabon102) follows from the linearity of the integral [5, Theorem 1.6.3]; and the equality in (LABEL:EqMartinezPabon105) follows from Theorem 1.
-On the other hand, for all $\mathcal{A}\in\mathscr{F}$
-$\displaystyle\int_{\mathcal{A}}f(x)\mathrm{d}P(x)$ $=$ $\displaystyle\int_{\mathcal{A}}\displaystyle\sum_{i=1}^{n}a_{i}\mathds{1}_{\mathcal{A}_{i}(x)}\mathrm{d}P(x)$ (6)
-$=$ $\displaystyle\sum_{i=1}^{n}\int_{\mathcal{A}}a_{i}\mathds{1}_{\mathcal{A}_{i}(x)}\mathrm{d}P(x)$ (7)
-$=$ $\sum_{i=1}^{n}a_{i}\int_{\mathcal{B}_{i}}\mathrm{d}P(x)=\sum_{i=1}^{n}a_{i}P(\mathcal{B}_{i}),$ (8)
-where the equality in (7) follows from the linearity of the integral [5, Theorem 1.6.3].
-Hence, from Theorem 1, and equalities (LABEL:EqMartinezPabon105) and (8), it follows that when $f$ is a simple function, the equality in (2) holds.
-This concludes the first part of the proof.
-The second part of the proof proceeds by considering the following observations: $\left(a\right)$ simple functions form a dense subset of the space of Borel measurable functions [5, Theorem 1.5.5(b)]; and $\left(b\right)$ the integral is a continuous map from that space [5, Theorem 1.6.2].
-Hence, from $\left(a\right)$ and $\left(b\right)$, it follows that (2) also holds for any Borel measurable function $f$.
+$=$ $\sum_{i=1}^{n}a_{i}\int_{\mathcal{B}_{i}}\frac{\mathrm{d}P}{\mathrm{d}Q}(x)\mathrm{d}Q(x)=\sum_{i=1}^{n}a_{i}P(\mathcal{B}_{i}),$ (4)
+where the equality in (4) follows Theorem 1 and the linearity of the integral [5, Theorem 1.6.3]. On the other hand, for all $\mathcal{A}\in\mathscr{F}$,
+$\int_{\mathcal{A}}f(x)\,\mathrm{d}P(x)$ $=$ $\sum_{i=1}^{m}a_{i}\int_{\mathcal{B}_{i}}\mathrm{d}P(x)\,=\,\sum_{i=1}^{m}a_{i}P(\mathcal{B}_{i}),$ (5)
+where the equality in (5) follows from the linearity of the integral [5, Theorem 1.6.3]. Hence, from Theorem 1 when $f$ is simple, the desired equality in (2) holds immediately.
+The proof proceeds by observing that: (a) simple functions are dense in the space of Borel measurable functions [5, Theorem 1.5.5(b)], and (b) the integral is a continuous map on that space [5, Theorem 1.6.2].
+From (a) and (b), it follows that (2) also holds for any Borel measurable function $f$.
 This completes the proof.
 ∎
-Another reputed folklore theorem, which is often referred to as the “proportional measures” theorem, establishes the explicit forms of the Radon-Nikodym derivatives between two measures, in which one is proportional to the other.
-Theorem 3 (Proportional Measures)
+Another reputed folklore theorem, which is often referred to as the “proportional measures” theorem, establishes the explicit form of the Radon-Nikodym derivative between two measures, in which one is proportional to the other.
+Theorem 3 (Proportional Measures).
 Let $P$ and $Q$ be two $\sigma$- finite measures on the measurable space $\left(\Omega,\mathscr{F}\right)$, such that for all $\mathcal{A}\in\mathscr{F}$
 $Q(\mathcal{A})=cP(\mathcal{A}),$ (6)
-with $c>0$. Then, for all $x\in\Omega$
-$\frac{\mathrm{d}P}{\mathrm{d}Q}(x)$ =$~{}\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{Q}\frac{1}{c},\mbox{ and }\frac{\mathrm{d}Q}{\mathrm{d}P}(x)$ =$\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{P}c.$ (10)
+with $c>0$. Then,
+$\frac{\mathrm{d}P}{\mathrm{d}Q}(x)\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{Q}\frac{1}{c},\quad\text{and}\quad\frac{\mathrm{d}Q}{\mathrm{d}P}(x)\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{P}c.$ (7)
 Proof:
-First, note that (9) implies that the measures $P$ and $Q$ are mutually absolutely continuous. Hence, from Theorem 1, it follows that for all $\mathcal{A}\in\mathscr{F}$,
-$P\left(\mathcal{A}\right)=\int_{\mathcal{A}}\mathrm{d}P(x)=\int_{\mathcal{A}}\frac{\mathrm{d}P}{\mathrm{d}Q}(x)\;\mathrm{d}Q(x).$ (11)
-On the other hand, the equality (9) also implies
-$P\left(\mathcal{A}\right)=\frac{1}{c}Q(\mathcal{A})=\int_{\mathcal{A}}\frac{1}{c}\mathrm{d}Q(x).$ (12)
-Hence, it follows directly from Theorem 1 that the Radon-Nikodym derivative $\frac{\mathrm{d}P}{\mathrm{d}Q}$ is unique almost surely with respect to $Q$.
-Thus, $\frac{\mathrm{d}P}{\mathrm{d}Q}(x)\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{Q}$$\frac{1}{c}$.
-Using similar arguments and the fact that $P$ and $Q$ are mutually absolutely continuous, it is verified that $\frac{\mathrm{d}Q}{\mathrm{d}P}(x)\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{P}$$c$.
+Note that from (6), $P$ and $Q$ are mutually absolutely continuous. Moreover, from Theorem 1, for all $\mathcal{A}\in\mathscr{F}$,
+$P(\mathcal{A})=\int_{\mathcal{A}}\mathrm{d}P(x)\;=\;\int_{\mathcal{A}}\frac{\mathrm{d}P}{\mathrm{d}Q}(x)\,\mathrm{d}Q(x),$ (8)
+and from (6), it follows that $P(\mathcal{A})=\frac{1}{c}\,Q(\mathcal{A})\;=\;\int_{\mathcal{A}}\frac{1}{c}\,\mathrm{d}Q(x)$.
+Hence, $\frac{\mathrm{d}P}{\mathrm{d}Q}(x)\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{Q}\frac{1}{c}$ follows directly from Theorem 1.
+A similar argument shows that $\frac{\mathrm{d}Q}{\mathrm{d}P}(x)\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{P}c$.
 ∎
-In the case in which $c=1$ in (10), measures $P$ and $Q$ are identical, thus, $\frac{\mathrm{d}P}{\mathrm{d}Q}(x)\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{Q}$ $\frac{\mathrm{d}Q}{\mathrm{d}P}(x)\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{P}$ $1$
-The following folklore theorem is often referred to as the “chain rule”.
-Theorem 4 (Chain Rule)
-Let $P$, $Q$, and $R$ be three measures on the measurable space $\left(\Omega,\mathscr{F}\right)$ such that ${P}\ll{Q}$; ${Q}\ll{R}$; and $Q$ and $R$ are $\sigma$-finite measures. Then,
+In the special case $c=1$ in (7), $P$ and $Q$ coincide, yielding $\frac{\mathrm{d}P}{\mathrm{d}Q}(x)\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{Q}$ $\frac{\mathrm{d}Q}{\mathrm{d}P}(x)\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{P}$ $1$
+The following folklore theorem is often referred to as the “chain rule” and a proof appears in [12, Proposition 3.9].
+Theorem 4 (Chain Rule).
+Let $P$ be a signed measure and let $Q$ and $R$ be two $\sigma$-finite measures, all on the same measurable space $\left(\Omega,\mathscr{F}\right)$, with ${P}\ll{Q}$ and ${Q}\ll{R}$. Then,
 $\frac{\mathrm{d}P}{\mathrm{d}R}(x)\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{R}\frac{\mathrm{d}P}{\mathrm{d}Q}(x)\frac{\mathrm{d}Q}{\mathrm{d}R}(x).$ (9)
-Proof:
-From the assumptions of the theorem, it follows that for all $\mathcal{A}\in\mathscr{F}$,
-$P(\mathcal{A})$ $=$ $\int_{\mathcal{A}}\mathrm{d}P(x)=\int_{\mathcal{A}}\frac{\mathrm{d}P}{\mathrm{d}Q}(x)\mathrm{d}Q(x)$ (14)
-$=$ $\int_{\mathcal{A}}\frac{\mathrm{d}P}{\mathrm{d}Q}(x)\frac{\mathrm{d}Q}{\mathrm{d}R}(x)\mathrm{d}R(x)$ (15)
-$=$ $\int_{\mathcal{A}}\frac{\mathrm{d}P}{\mathrm{d}R}(x)\mathrm{d}R(x),$ (16)
-where the second equality in (14) follows from Theorem 1; and the equality in (15) follows from Theorem 2 .
-The equality in (16) holds from Theorem 1 and by noticing that $P~{}\ll~{}R$.
-Therefore, the equalities in (15) and (16) together with Theorem 1 imply (13), which completes the proof.
-∎
-The following folklore theorem shows the connection between the Radon-Nikodym derivative and its multiplicative inverse.
-Theorem 5 (Multiplicative Inverse)
-Let $P$ and $Q$ be two mutually absolutely continuous measures on the measurable space $\left(\Omega,\mathscr{F}\right)$; and assume that for all $x\in\Omega$, $\frac{\mathrm{d}Q}{\mathrm{d}P}(x)>0$.
+The following folklore theorem is a consequence of Theorem 4 and shows the connection between the Radon-Nikodym derivative and its multiplicative inverse.
+This result also appears in [12, Corollary 3.10].
+Theorem 5 (Multiplicative Inverse).
+Let $P$ and $Q$ be two mutually absolutely continuous $\sigma$-finite measures on the measurable space $\left(\Omega,\mathscr{F}\right)$; and assume that for all $x\in\Omega$, $\frac{\mathrm{d}Q}{\mathrm{d}P}(x)>0$.
 Then,
 $\frac{\mathrm{d}P}{\mathrm{d}Q}(x)$ =$\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{Q}$ $\left(\frac{\mathrm{d}Q}{\mathrm{d}P}(x)\right)^{-1}.$ (10)
+The subsequent folklore theorem establishes the linearity of the Radon-Nikodym derivative. A similar statement is presented in [12, Proposition 3.11] but left without proof.
+Theorem 6 (Linearity).
+Let $P$ be a $\sigma$-finite measure on $\left(\Omega,\mathscr{F}\right)$ and let also $Q_{1},Q_{2},\ldots,Q_{n}$ be signed measures on $\left(\Omega,\mathscr{F}\right)$ absolutely continuous with respect to $P$.
+Let $c_{1},c_{2},\ldots,c_{n}$ be positive reals; and let $S$ be a signed measure on $\left(\Omega,\mathscr{F}\right)$ such that for all $\mathcal{A}\in\mathscr{F}$, $S(\mathcal{A})=\sum_{t=1}^{n}c_{t}Q_{t}(\mathcal{A})$.
+Then,
+$\frac{\mathrm{d}S}{\mathrm{d}P}(x)\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{P}\sum_{t=1}^{n}c_{t}\frac{\mathrm{d}Q_{t}}{\mathrm{d}P}(x).$ (11)
 Proof:
-From Theorem 4, it follows that
-$\frac{\mathrm{d}P}{\mathrm{d}Q}(x)\frac{\mathrm{d}Q}{\mathrm{d}P}(x)\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{Q}\frac{\mathrm{d}Q}{\mathrm{d}Q}(x)\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{Q}1,$ (18)
-where the last equality follows from Theorem 3, with $c~{}=~{}1$. This completes the proof. ∎
-The subsequent folklore theorem establishes the linearity of the Radon-Nikodym derivative.
-Theorem 6 (Linearity)
-Let $P$ be a $\sigma$-finite measure on $\left(\Omega,\mathscr{F}\right)$ and let also $Q_{1},Q_{2},\ldots,Q_{n}$ be finite measures on $\left(\Omega,\mathscr{F}\right)$ absolutely continuous with respect to $P$.
-Let $c_{1},c_{2},\ldots,c_{n}$ be positive reals; and let $S$ be a finite measure on $\left(\Omega,\mathscr{F}\right)$ such that for all $\mathcal{A}\in\mathscr{F}$, $S(\mathcal{A})=\sum_{t=1}^{n}c_{t}Q_{t}(\mathcal{A})$.
-Then,
-$\frac{\mathrm{d}S}{\mathrm{d}P}(x)$ =$\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{P}$ $\sum_{t=1}^{n}c_{t}\frac{\mathrm{d}Q_{t}}{\mathrm{d}P}(x).$ (19)
-Proof:
-The proof starts by noticing that, from the assumptions of the theorem, it holds that ${S}\ll{P}$. Hence, for all $\mathcal{A}\in\mathscr{F}$, it holds that
-$\int_{\mathcal{A}}\frac{\mathrm{d}S}{\mathrm{d}P}(x)\mathrm{d}P(x)=\int_{\mathcal{A}}\mathrm{d}S(x)=\sum_{t=1}^{n}c_{t}Q_{t}(\mathcal{A})$ (20)
-$=$ $\sum_{t=1}^{n}\int_{\mathcal{A}}c_{t}\mathrm{d}Q_{t}(x)=\sum_{t=1}^{n}\int_{\mathcal{A}}c_{t}\frac{\mathrm{d}Q_{t}}{\mathrm{d}P}(x)\mathrm{d}P(x)$ (21)
-$=$ $\int_{\mathcal{A}}\sum_{t=1}^{n}c_{t}\frac{\mathrm{d}Q_{t}}{\mathrm{d}P}(x)\mathrm{d}P(x),$ (22)
-where the first equality in (20) and the last equality in (21) follow from Theorem 2; and the equality (22) follows from the additivity property of the integral [5, Corollary 1.6.4].
-The proof ends by using Theorem 1, which implies the equality in (19) from (22).
+From the assumptions of the theorem, it holds that ${S}\ll{P}$, thus for all $\mathcal{A}\in\mathscr{F}$,
+$\int_{\mathcal{A}}$ $\frac{\mathrm{d}S}{\mathrm{d}P}(x)\mathrm{d}P(x)=\int_{\mathcal{A}}\mathrm{d}S(x)=\sum_{t=1}^{n}\int_{\mathcal{A}}c_{t}\mathrm{d}Q_{t}(x)$ (12)
+$=$ $\sum_{t=1}^{n}\int_{\mathcal{A}}c_{t}\frac{\mathrm{d}Q_{t}}{\mathrm{d}P}(x)\mathrm{d}P(x)=\int_{\mathcal{A}}\sum_{t=1}^{n}c_{t}\frac{\mathrm{d}Q_{t}}{\mathrm{d}P}(x)\mathrm{d}P(x),$ (13)
+where the first equality in (12) and the first equality in (13) follow from Theorem 2; and the last equality in (13) follows from the additivity property of the integral [5, Corollary 1.6.4].
+The proof ends by using Theorem 1 yielding (11).
 ∎
 The following folklore theorem establishes the continuity of the Radon-Nikodym derivative.
-Theorem 7 (Continuity)
-Let $P$ be a $\sigma$-finite measure on $\left(\Omega,\mathscr{F}\right)$, and let $Q_{1},Q_{2},\cdots$ be an infinite sequence of $\sigma$-finite measures on $\left(\Omega,\mathscr{F}\right)$, converging to a measure $Q$.
+Theorem 7 (Continuity).
+Let $P$ be a $\sigma$-finite measure on $\left(\Omega,\mathscr{F}\right)$, and let $Q_{1},Q_{2},\cdots$ be an infinite sequence of signed measures on $\left(\Omega,\mathscr{F}\right)$, converging to a signed measure $Q$.
 Suppose that for all $n\in\mathbb{N}$, ${Q_{n}}\ll{P}$.
@@
 Proof:
-From the assumptions of the theorem, for all $\mathcal{A}\in\mathscr{F}$, it holds that
-$Q(\mathcal{A})$ $=$ $\lim_{n\rightarrow\infty}Q_{n}(\mathcal{A})$ (24)
-$=$ $\lim_{n\rightarrow\infty}\int_{\mathcal{A}}\frac{\mathrm{d}Q_{n}}{\mathrm{d}P}(x)\mathrm{d}P(x)$ (25)
-$=$ $\int_{\mathcal{A}}\lim_{n\to\infty}\frac{\mathrm{d}Q_{n}}{\mathrm{d}P}(x)\mathrm{d}P(x),$ (26)
-where the equality in (25) follows from Theorem 2 ; and the equality in (26) follows from [5, Theorem 1.6.2].
-The equality in (26) implies that ${Q}\ll{P}$.
-Hence, for all $\mathcal{A}\in\mathscr{F}$, it holds that
+From the assumptions of the theorem and Theorem 2, for all $\mathcal{A}\in\mathscr{F}$, it holds that
+$Q(\mathcal{A})$ $=$ $\lim_{n\to\infty}Q_{n}(\mathcal{A})\;=\;\int_{\mathcal{A}}\lim_{n\to\infty}\frac{\mathrm{d}Q_{n}}{\mathrm{d}P}(x)\,\mathrm{d}P(x),$ (15)
+where the last equality in (15) follows from Theorem 2 and [5, Theorem 1.6.2]. Hence $Q\ll P$ and for all $\mathcal{A}\in\mathscr{F}$, it holds that
 $Q(\mathcal{A})\;=\;\int_{\mathcal{A}}\frac{\mathrm{d}Q}{\mathrm{d}P}(x)\,\mathrm{d}P(x).$ (16)
-Therefore, the equalities in (26) and (27) jointly with Theorem 1 imply equation (23), which completes the proof. ∎
-The ensuing folklore theorem establishes the relation between the Radon-Nikodym derivative of a product measure with respect to its component measures.
-Theorem 8 (Product of Measures)
-For all $i\in\{1,2\}$, let $P_{i}$ and $Q_{i}$ be a finite and a $\sigma$-finite measure on $\left(\Omega_{i},\mathscr{F}_{i}\right)$, respectively; with ${P_{i}}\ll{Q_{i}}$.
+Therefore, the equality in (16) and Theorem 1 yields (14). ∎
+The ensuing folklore theorem relates the Radon–Nikodym derivative of a product measure to those of its component measures. This result also appears in [12, Exercise 3.12].
+Theorem 8 (Product of Measures).
+For all $i\in\{1,2\}$, let $P_{i}$ and $Q_{i}$ be a signed and a $\sigma$-finite measure on $\left(\Omega_{i},\mathscr{F}_{i}\right)$, respectively; with ${P_{i}}\ll{Q_{i}}$.
 Let also $P_{1}P_{2}$ and $Q_{1}Q_{2}$ be the product measures on $\left(\Omega_{1}\times\Omega_{2},\mathscr{F}_{1}\times\mathscr{F}_{2}\right)$ formed by $P_{1}$ and $P_{2}$; and $Q_{1}$ and $Q_{2}$, respectively.
@@
 Proof:
-From the assumptions of the theorem, for all $\mathcal{A}\in\left(\Omega_{1}\times\Omega_{2}\right)$,
-$P_{1}P_{2}\left(\mathcal{A}\right)$ $=$ $\int_{\mathcal{A}}\mathrm{d}P_{1}P_{2}\left(x_{1},x_{2}\right)$ (29)
-$=$ $\int\int_{\mathcal{A}_{x_{2}}}\mathrm{d}P_{1}\left(x_{1}\right)\mathrm{d}P_{2}\left(x_{2}\right)$ (30)
-$=$ $\int\int_{\mathcal{A}_{x_{2}}}\dfrac{\mathrm{d}P_{1}\left(x_{1}\right)}{\mathrm{d}Q_{1}}\mathrm{d}Q_{1}\left(x_{1}\right)\mathrm{d}P_{2}\left(x_{2}\right)$ (31)
-$=$ $\int\int_{\mathcal{A}_{x_{2}}}\dfrac{\mathrm{d}P_{1}\left(x_{1}\right)}{\mathrm{d}Q_{1}}\dfrac{\mathrm{d}P_{2}\left(x_{2}\right)}{\mathrm{d}Q_{2}}\mathrm{d}Q_{1}\left(x_{1}\right)\mathrm{d}Q_{2}\left(x_{2}\right)$ (32)
+By definition of product measures, for all $\mathcal{A}\subseteq\Omega_{1}\times\Omega_{2}$,
+$P_{1}P_{2}\left(\mathcal{A}\right)$ $=$ $\int_{\mathcal{A}}\mathrm{d}P_{1}P_{2}\left(x_{1},x_{2}\right)=\iint_{\mathcal{A}_{x_{2}}}\mathrm{d}P_{1}\left(x_{1}\right)\mathrm{d}P_{2}\left(x_{2}\right)$ (18)
+$=$ $\iint_{\mathcal{A}_{x_{2}}}\dfrac{\mathrm{d}P_{1}\left(x_{1}\right)}{\mathrm{d}Q_{1}}\mathrm{d}Q_{1}\left(x_{1}\right)\mathrm{d}P_{2}\left(x_{2}\right)$ (19)
+$=$ $\iint_{\mathcal{A}_{x_{2}}}\dfrac{\mathrm{d}P_{1}\left(x_{1}\right)}{\mathrm{d}Q_{1}}\dfrac{\mathrm{d}P_{2}\left(x_{2}\right)}{\mathrm{d}Q_{2}}\mathrm{d}Q_{1}\left(x_{1}\right)\mathrm{d}Q_{2}\left(x_{2}\right)$ (20)
 $=$ $\int_{\mathcal{A}}\dfrac{\mathrm{d}P_{1}}{\mathrm{d}Q_{1}}\left(x_{1}\right)\dfrac{\mathrm{d}P_{2}}{\mathrm{d}Q_{2}}\left(x_{2}\right)\mathrm{d}Q_{1}Q_{2}\left(x_{1},x_{2}\right),$ (21)
-where $\mathcal{A}_{x_{2}}$ is the section of the set $\mathcal{A}$ determined by $x_{2}$, namely, $\mathcal{A}_{x_{2}}\triangleq\left\{x_{1}\in\Omega_{1}:(x_{1},x_{2})\in\mathcal{A}\right\}$; the equality in (29) arises from the definition of $P_{1}P_{2}$ as the product of $P_{1}$ and $P_{2}$; the equality in (31) is a direct consequence of Theorem 2; the equality in (32) follows from Theorem 1; and finally, the equality in (33) is due to the construction of $Q_{1}Q_{2}$ as the product measure of $Q_{1}$ and $Q_{2}$.
-The proof follows by observing that from the equality in (33), it holds that ${P_{1}P_{2}}\ll{Q_{1}Q_{2}}$. Thus, for all $\mathcal{A}~{}\in~{}\mathscr{F}_{1}~{}\times~{}\mathscr{F}_{2}$,
+where $\mathcal{A}_{x_{2}}$ is the section of the set $\mathcal{A}$ determined by $x_{2}$, namely, $\mathcal{A}_{x_{2}}=\{x_{1}:(x_{1},x_{2})\in\mathcal{A}\}$; the equality in (18) arises from the definition of $P_{1}P_{2}$ as the product of $P_{1}$ and $P_{2}$; the equalities in (19) and (20) are direct consequences of Theorem 2; and finally, the equality in (21) is due to the construction of $Q_{1}Q_{2}$ as the product measure of $Q_{1}$ and $Q_{2}$.
+From the equality in (21), it holds that ${P_{1}P_{2}}\ll{Q_{1}Q_{2}}$.
+Thus, for all $\mathcal{A}\leavevmode\nobreak\ \in\leavevmode\nobreak\ \mathscr{F}_{1}\leavevmode\nobreak\ \times\leavevmode\nobreak\ \mathscr{F}_{2}$,
 $P_{1}P_{2}\left(\mathcal{A}\right)=\int_{\mathcal{A}}\dfrac{\mathrm{d}P_{1}P_{2}}{\mathrm{d}Q_{1}Q_{2}}\left(x_{1},x_{2}\right)\mathrm{d}Q_{1}Q_{2}\left(x_{1},x_{2}\right).$ (22)
-The equalities in (33) and (34), together with Theorem 1, imply the equality in (28), which completes the proof. ∎
+From Theorem 1, the equality in (17) follows, completing the proof. ∎
 IV Advanced folklore theorems
-This section requires some additional notation.
-In particular, denote by $\triangle\left(\mathcal{X},\mathscr{F}_{\mathcal{X}}\right)$, or simply $\triangle\left(\mathcal{X}\right)$, the set of all probability measures on the measurable space $\left(\mathcal{X},\mathscr{F}_{\mathcal{X}}\right)$, where $\mathscr{F}_{\mathcal{X}}$ is a $\sigma$-algebra on $\mathcal{X}$.
+This section focuses exclusively on probability measures with the aim of establishing closer ties to information measures, which requires some additional notation.
+In particular, denote by $\triangle\bigl{(}\mathcal{X},\mathscr{F}_{\mathcal{X}}\bigr{)}$, or simply $\triangle\bigl{(}\mathcal{X}\bigr{)}$, the set of all probability measures on the measurable space $\bigl{(}\mathcal{X},\mathscr{F}_{\mathcal{X}}\bigr{)}$, where $\mathscr{F}_{\mathcal{X}}$ is a $\sigma$-algebra on $\mathcal{X}$.
 Using this notation, conditional probability measures can be defined as follows.
-Definition 1 (Conditional Probability)
-A family $P_{Y|X}~{}\triangleq~{}(P_{Y|X=x})_{x\in\mathcal{X}}$ of elements of $\triangle(\mathcal{Y},\mathscr{F}_{\mathcal{Y}})$ indexed by $\mathcal{X}$ is said to be a conditional probability measure if, for all sets $\mathcal{A}\in\mathscr{F}_{\mathcal{Y}}$, the map
-$\mathcal{X}$$\to$ $[0,1]$ (35)
-$x$ $\mapsto$ $P_{Y|X=x}(\mathcal{A})$ (36)
-is Borel measurable. The set of such conditional probability measures is denoted by $\triangle(\mathcal{Y}|\mathcal{X})$.
-A conditional probability $P_{Y|X}\in\triangle(\mathcal{Y}|\mathcal{X})$ and a probability measure $P_{X}\in\triangle(\mathcal{X})$ determine two unique probability measures in $\triangle(\mathcal{X}\times\mathcal{Y})$ and $\triangle(\mathcal{Y}\times\mathcal{X})$, respectively.
+Definition 1 (Conditional Probability).
+A family $P_{Y|X}\,\triangleq\,(P_{Y|X=x})_{x\in\mathcal{X}}$ of elements of $\triangle\bigl{(}\mathcal{Y},\mathscr{F}_{\mathcal{Y}}\bigr{)}$ indexed by $\mathcal{X}$ is said to be a conditional probability measure if, for all sets $\mathcal{B}\in\mathscr{F}_{\mathcal{Y}}$, the map
+[…]
+is Borel measurable. The set of such conditional probability measures is denoted by $\triangle(\mathcal{Y}\!\mid\!\mathcal{X})$.
+A conditional probability $P_{Y|X}\in\triangle(\mathcal{Y}\mid\mathcal{X})$ and a probability measure $P_{X}\in\triangle(\mathcal{X})$ jointly determine two unique probability measures in $\triangle(\mathcal{X}\times\mathcal{Y})$ and $\triangle(\mathcal{Y}\times\mathcal{X})$, respectively.
 These probability measures are denoted by $P_{XY}$ and $P_{YX}$, respectively, and for all sets $\mathcal{A}\in\mathscr{F}_{\mathcal{X}}\times\mathscr{F}_{\mathcal{Y}}$, it follows that
-$P_{XY}\left(\mathcal{A}\right)$ $=$ $\int P_{Y|X=x}\left(\mathcal{A}_{x}\right)\mathrm{d}P_{X}\left(x\right),$ (37)
+$P_{XY}\bigl{(}\mathcal{A}\bigr{)}$ $=$ $\int P_{Y|X=x}\bigl{(}\mathcal{A}_{x}\bigr{)}\,\mathrm{d}P_{X}(x),$ (24)
 where $\mathcal{A}_{x}$ is the section of the set $\mathcal{A}$ determined by $x$, namely,
-$\mathcal{A}_{x}\triangleq\left\{y\in\mathcal{Y}:(x,y)\in\mathcal{A}\right\}.$ (38)
+$\mathcal{A}_{x}$$\triangleq$ $\bigl{\{}y\in\mathcal{Y}:(x,y)\in\mathcal{A}\bigr{\}}.$ (25)
 Alternatively, for all sets $\mathcal{B}\in\mathscr{F}_{\mathcal{Y}}\times\mathscr{F}_{\mathcal{X}}$, it follows that
-$P_{YX}\left(\mathcal{B}\right)$ $=$ $\int P_{Y|X=x}\left(\mathcal{B}_{x}\right)\mathrm{d}P_{X}\left(x\right),$ (39)
+$P_{YX}\bigl{(}\mathcal{B}\bigr{)}=\int P_{Y|X=x}\bigl{(}\mathcal{B}_{x}\bigr{)}\,\mathrm{d}P_{X}(x),$ (26)
 where $\mathcal{B}_{x}$ is the section of the set $\mathcal{B}$ determined by $x$.
-For all sets $\mathcal{A}\in\mathscr{F}_{\mathcal{X}}\times\mathscr{F}_{\mathcal{Y}}$, let the set $\hat{\mathcal{A}}\in\mathscr{F}_{\mathcal{Y}}\times\mathscr{F}_{\mathcal{X}}$ be such that
-$\hat{\mathcal{A}}$ $=$ $\left\{(y,x)\in\mathcal{Y}\times\mathcal{X}:(x,y)\in\mathcal{A}\right\}.$ (40)
+Note that $P_{XY}$ is a measure in $\triangle\left(\mathcal{X}\times\mathcal{Y}\right)$ whereas $P_{YX}$ is a measure in $\triangle\left(\mathcal{Y}\times\mathcal{X}\right)$.
+Consequently, obtaining $P_{XY}$ from $P_{YX}$ requires more than simply renaming $X$ and $Y$; one must also properly redefine the corresponding measurable spaces.
+In particular, given a set $\mathcal{A}\in\mathscr{F}_{\mathcal{X}}\times\mathscr{F}_{\mathcal{Y}}$, let the set $\widehat{\mathcal{A}}\in\mathscr{F}_{\mathcal{Y}}\times\mathscr{F}_{\mathcal{X}}$ be such that
+$\widehat{\mathcal{A}}=\bigl{\{}(y,x)\in\mathcal{Y}\times\mathcal{X}:(x,y)\in\mathcal{A}\bigr{\}}.$ (27)
 Then, from (24) and (26), it holds that
-$P_{XY}\left(\mathcal{A}\right)$ $=$ $P_{YX}\left(\hat{\mathcal{A}}\right).$ (41)
-Using this notation, the notion of marginal probability measures can be introduced as follows.
-Definition 2 (Marginals)
-Given two joint probability measures $P_{XY}\in\triangle\left(\mathcal{X}\times\mathcal{Y}\right)$ and $P_{YX}\in\triangle\left(\mathcal{Y}\times\mathcal{X}\right)$, satisfying (41), the marginal probability measures in $\triangle\left(\mathcal{X}\right)$ and $\triangle\left(\mathcal{Y}\right)$, denoted by $P_{X}$ and $P_{Y}$, respectively satisfy for all sets $\mathcal{A}\in\mathscr{F}_{\mathcal{X}}$ and for all sets $\mathcal{B}\in\mathscr{F}_{\mathcal{Y}}$,
-$P_{X}\left(\mathcal{A}\right)$ $\triangleq P_{XY}\left(\mathcal{A}\times\mathcal{Y}\right)=$ $P_{YX}\left(\mathcal{Y}\times\mathcal{A}\right);\mbox{ and }$ (42)
-$P_{Y}\left(\mathcal{B}\right)$ $\triangleq P_{XY}\left(\mathcal{X}\times\mathcal{B}\right)=$ $P_{YX}\left(\mathcal{B}\times\mathcal{X}\right).$ (43)
-From the total probability theorem [5, Theorem 4.5.2], it follows that for all $\mathcal{A}\in\mathscr{F}_{\mathcal{Y}}$,
-$P_{Y}(\mathcal{A})=\int\int_{\mathcal{A}}\mathrm{d}P_{Y|X=x}(y)\mathrm{d}P_{X}(x);$ (44)
-and for all $\mathcal{B}\in\mathscr{F}_{\mathcal{X}}$,
-$P_{X}(\mathcal{B})=\int\int_{\mathcal{B}}\mathrm{d}P_{X|Y=y}(x)\mathrm{d}P_{Y}(y).$ (45)
-The joint probability measures $P_{XY}$ and $P_{YX}$ can be described via the conditional probability measure $P_{Y|X}$ and the probability measure $P_{X}$ as in (37) and in (39); or via the conditional probability measure $P_{X|Y}\in\triangle\left(\mathcal{X}|\mathcal{Y}\right)$ and the marginal probability measure $P_{Y}\in\triangle\left(\mathcal{Y}\right)$.
+$P_{XY}\bigl{(}\mathcal{A}\bigr{)}=P_{YX}\bigl{(}\widehat{\mathcal{A}}\bigr{)}.$ (28)
+Using this notation, a marginal probability measure is defined as follows.
+Definition 2 (Marginal).
+Given two joint probability measures $P_{XY}\in\triangle\left(\mathcal{X}\times\mathcal{Y}\right)$ and $P_{YX}\in\triangle\left(\mathcal{Y}\times\mathcal{X}\right)$, satisfying (28), the marginal probability measures in $\triangle\left(\mathcal{X}\right)$ and $\triangle\left(\mathcal{Y}\right)$, denoted by $P_{X}$ and $P_{Y}$, respectively, satisfy for all sets $\mathcal{A}\in\mathscr{F}_{\mathcal{X}}$
+$P_{X}\left(\mathcal{A}\right)\triangleq P_{XY}\left(\mathcal{A}\times\mathcal{Y}\right)=P_{YX}\left(\mathcal{Y}\times\mathcal{A}\right);$ (29)
+and for all sets $\mathcal{B}\in\mathscr{F}_{\mathcal{Y}}$,
+$P_{Y}\left(\mathcal{B}\right)\triangleq P_{YX}\left(\mathcal{B}\times\mathcal{X}\right)=P_{XY}\left(\mathcal{X}\times\mathcal{B}\right).$ (30)
+From the total probability theorem [5, Theorem 4.5.2], it follows that for all $\mathcal{B}\in\mathscr{F}_{\mathcal{Y}}$,
+$P_{Y}(\mathcal{B})=\iint_{\mathcal{B}}\mathrm{d}P_{Y|X=x}(y)\mathrm{d}P_{X}(x);$ (31)
+and for all $\mathcal{A}\in\mathscr{F}_{\mathcal{X}}$,
+$P_{X}(\mathcal{A})=\iint_{\mathcal{A}}\mathrm{d}P_{X|Y=y}(x)\mathrm{d}P_{Y}(y).$ (32)
+In a nutshell, the joint probability measures $P_{XY}$ and $P_{YX}$ can be described via the conditional probability measure $P_{Y|X}$ and the probability measure $P_{X}$ as in (24) and in (26); or via the conditional probability measure $P_{X|Y}\in\triangle\left(\mathcal{X}|\mathcal{Y}\right)$ and the marginal probability measure $P_{Y}\in\triangle\left(\mathcal{Y}\right)$.
 More specifically, for all sets $\mathcal{A}\in\mathscr{F}_{\mathcal{X}}\times\mathscr{F}_{\mathcal{Y}}$, it follows that
@@
 Within this context, the following folklore theorem highlights a property of conditional measures, which is reminiscent of the unit measure axiom in probability theory.
-Theorem 9 (Unit Measure)
+Theorem 9 (Unit Measure).
 Consider the conditional probability measures $P_{Y|X}\in\triangle\left(\mathcal{Y}|\mathcal{X}\right)$ and $P_{X|Y}\in\triangle\left(\mathcal{X}|\mathcal{Y}\right)$; the probability measures $P_{Y}\in\triangle\left(\mathcal{Y}\right)$ and $P_{X}\in\triangle\left(\mathcal{X}\right)$ that satisfy (31) and (32).
@@
 For all $\mathcal{A}\in\mathscr{F}_{\mathcal{Y}}$, from (31), it holds that
-$P_{Y}(\mathcal{A})$ $=$ $\int\int_{\mathcal{A}}\mathrm{d}P_{Y|X=x}(y)\mathrm{d}P_{X}(x)$ (50)
-$=$ $\int\int_{\mathcal{A}}\frac{\mathrm{d}P_{Y|X=x}}{\mathrm{d}P_{Y}}(y)\mathrm{d}P_{Y}(y)\mathrm{d}P_{X}(x)$ (51)
+$P_{Y}(\mathcal{A})$ $=$ $\iint_{\mathcal{A}}\mathrm{d}P_{Y|X=x}(y)\mathrm{d}P_{X}(x)$ (37)
+$=$ $\iint_{\mathcal{A}}\frac{\mathrm{d}P_{Y|X=x}}{\mathrm{d}P_{Y}}(y)\mathrm{d}P_{Y}(y)\mathrm{d}P_{X}(x)$ (38)
 $=$ $\int_{\mathcal{A}}\int\frac{\mathrm{d}P_{Y|X=x}}{\mathrm{d}P_{Y}}(y)\mathrm{d}P_{X}(x)\mathrm{d}P_{Y}(y),$ (39)
@@
 ∎
-The following folklore theorem is reminiscent of the Bayes rule.
-Theorem 10 (Bayes-like rule)
+The following folklore theorems are reminiscent of the Bayes rule.
+Theorem 10 (Bayes-like rule).
 Consider the conditional probability measures $P_{Y|X}$ and $P_{X|Y}$; the probability measures $P_{Y}$ and $P_{X}$ that satisfy (31) and (32); and the joint probability measures $P_{YX}$ and $P_{XY}$ in (26) and (33) respectively.
 Let also $P_{X}P_{Y}\in\triangle\left(\mathcal{X}\times\mathcal{Y},\mathscr{F}_{\mathcal{X}}\times\mathscr{F}_{\mathcal{Y}}\right)$ and $P_{Y}P_{X}\in\triangle\left(\mathcal{Y}\times\mathcal{X},\mathscr{F}_{\mathcal{Y}}\times\mathscr{F}_{\mathcal{X}}\right)$ be the measures formed by the product of the marginals $P_{X}$ and $P_{Y}$.
-Assume that:
-  • $\left(a\right)$
-    For all $x\in\mathcal{X}$${P_{Y|X=x}}\ll{P_{Y}}$; and
-  • $\left(b\right)$
-    For all $y\in\mathcal{Y}$${P_{X|Y=y}}\ll{P_{X}}$.
+Assume that: $\left(a\right)$ For all $x\in\mathcal{X}$${P_{Y|X=x}}\ll{P_{Y}}$; and $\left(b\right)$ For all $y\in\mathcal{Y}$${P_{X|Y=y}}\ll{P_{X}}$.
 Then,
 $\frac{\mathrm{d}P_{XY}}{\mathrm{d}P_{X}P_{Y}}\left(x,y\right)$ =$\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{P_{X}P_{Y}}$ $\frac{\mathrm{d}P_{X|Y=y}}{\mathrm{d}P_{X}}\left(x\right)$ (40)
-=$\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{P_{X}P_{Y}}$ $\frac{\mathrm{d}P_{Y|X=x}}{\mathrm{d}P_{Y}}\left(y\right)$ (54)
-=$\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{P_{X}P_{Y}}$ $\frac{\mathrm{d}P_{YX}}{\mathrm{d}P_{Y}P_{X}}\left(y,x\right).$ (55)
+=$\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{P_{X}P_{Y}}\frac{\mathrm{d}P_{Y|X=x}}{\mathrm{d}P_{Y}}\left(y\right)$ =$\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{P_{X}P_{Y}}$ $\frac{\mathrm{d}P_{YX}}{\mathrm{d}P_{Y}P_{X}}\left(y,x\right).$ (41)
 Proof:
-Note that assumptions $\left(a\right)$ and $\left(b\right)$ are sufficient for the Radon-Nikodym derivatives of $P_{XY}$ with respect to $P_{X}P_{Y}$ and $P_{YX}$ with respect to $P_{Y}P_{X}$ to exist.
-Hence, it follows that for all sets $\mathcal{A}~{}\in~{}\mathscr{F}_{\mathcal{X}}\times\mathscr{F}_{\mathcal{Y}}$,
+Note that assumptions $\left(a\right)$ and $\left(b\right)$ are sufficient for the Radon-Nikodym derivatives of $P_{XY}$ with respect to $P_{X}P_{Y}$; and $P_{YX}$ with respect to $P_{Y}P_{X}$ to exist.
+Hence, it follows that for all sets $\mathcal{A}\leavevmode\nobreak\ \in\leavevmode\nobreak\ \mathscr{F}_{\mathcal{X}}\times\mathscr{F}_{\mathcal{Y}}$,
 $P_{XY}\left(\mathcal{A}\right)$ $=$ $\int_{\mathcal{A}}\frac{\mathrm{d}P_{XY}}{\mathrm{d}P_{X}P_{Y}}(x,y)\mathrm{d}P_{X}P_{Y}(x,y),$ (42)
 which follows from Theorem 2. Note also that from (33), it follows that
-$P_{XY}\left(\mathcal{A}\right)$ $=$ $\int\int_{\mathcal{A}_{y}}\mathrm{d}P_{X|Y=y}\left(x\right)\mathrm{d}P_{Y}\left(y\right)$ (57)
-$=$ $\int\int_{\mathcal{A}_{y}}\frac{\mathrm{d}P_{X|Y=y}}{\mathrm{d}P_{X}}\left(x\right)\mathrm{d}P_{X}\left(x\right)\mathrm{d}P_{Y}\left(y\right)$ (58)
-$=$ $\int\int\mathds{1}_{\mathcal{A}_{y}}(x)\frac{\mathrm{d}P_{X|Y=y}}{\mathrm{d}P_{X}}\left(x\right)\mathrm{d}P_{X}\left(x\right)\mathrm{d}P_{Y}\left(y\right)$ (59)
+$P_{XY}\left(\mathcal{A}\right)$ $=$ $\iint_{\mathcal{A}_{y}}\mathrm{d}P_{X|Y=y}\left(x\right)\mathrm{d}P_{Y}\left(y\right)$ (43)
+$=$ $\iint_{\mathcal{A}_{y}}\frac{\mathrm{d}P_{X|Y=y}}{\mathrm{d}P_{X}}\left(x\right)\mathrm{d}P_{X}\left(x\right)\mathrm{d}P_{Y}\left(y\right)$ (44)
+$=$ $\iint\mathds{1}_{\mathcal{A}_{y}}(x)\frac{\mathrm{d}P_{X|Y=y}}{\mathrm{d}P_{X}}\left(x\right)\mathrm{d}P_{X}\left(x\right)\mathrm{d}P_{Y}\left(y\right)$ (45)
 $=$ $\int\mathds{1}_{\mathcal{A}}(x,y)\frac{\mathrm{d}P_{X|Y=y}}{\mathrm{d}P_{X}}\left(x\right)\mathrm{d}P_{X}P_{Y}\left(x,y\right)$ (46)
@@
 where, the set $\mathcal{A}_{y}$ is defined in (34). Moreover, the equality in (44) follows from Assumption $\left(b\right)$ and Theorem 1. Similarly, from (24), it follows that
-$P_{XY}\left(\mathcal{A}\right)$ $=$ $\int\int_{\mathcal{A}_{x}}\mathrm{d}P_{Y|X=x}\left(y\right)\mathrm{d}P_{X}\left(x\right)$ (62)
-$=$ $\int\int_{\mathcal{A}_{x}}\frac{\mathrm{d}P_{Y|X=x}}{\mathrm{d}P_{Y}}\left(y\right)\mathrm{d}P_{Y}\left(y\right)\mathrm{d}P_{X}\left(x\right)$ (63)
-$=$ $\int\int\mathds{1}_{\mathcal{A}_{x}}(y)\frac{\mathrm{d}P_{Y|X=x}}{\mathrm{d}P_{Y}}\left(y\right)\mathrm{d}P_{Y}\left(y\right)\mathrm{d}P_{X}\left(x\right)$ (64)
-$=$ $\int\int\mathds{1}_{\mathcal{A}_{y}}(x)\frac{\mathrm{d}P_{Y|X=x}}{\mathrm{d}P_{Y}}\left(y\right)\mathrm{d}P_{X}\left(x\right)\mathrm{d}P_{Y}\left(y\right)$ (65)
+$P_{XY}\left(\mathcal{A}\right)$ $=$ $\iint_{\mathcal{A}_{x}}\mathrm{d}P_{Y|X=x}\left(y\right)\mathrm{d}P_{X}\left(x\right)$ (48)
+$=$ $\iint_{\mathcal{A}_{x}}\frac{\mathrm{d}P_{Y|X=x}}{\mathrm{d}P_{Y}}\left(y\right)\mathrm{d}P_{Y}\left(y\right)\mathrm{d}P_{X}\left(x\right)$ (49)
+$=$ $\iint\mathds{1}_{\mathcal{A}_{x}}(y)\frac{\mathrm{d}P_{Y|X=x}}{\mathrm{d}P_{Y}}\left(y\right)\mathrm{d}P_{Y}\left(y\right)\mathrm{d}P_{X}\left(x\right)$ (50)
+$=$ $\iint\mathds{1}_{\mathcal{A}_{y}}(x)\frac{\mathrm{d}P_{Y|X=x}}{\mathrm{d}P_{Y}}\left(y\right)\mathrm{d}P_{X}\left(x\right)\mathrm{d}P_{Y}\left(y\right)$ (51)
 $=$ $\int\mathds{1}_{\mathcal{A}}(x,y)\frac{\mathrm{d}P_{Y|X=x}}{\mathrm{d}P_{Y}}\left(y\right)\mathrm{d}P_{X}P_{Y}\left(x,y\right)$ (52)
@@
 Finally, from (28), it follows that
-$P_{XY}\left(\mathcal{A}\right)=\int_{\hat{\mathcal{A}}}\mathrm{d}P_{YX}\left(y,x\right)$ (68)
-$=$ $\int_{\hat{\mathcal{A}}}\frac{\mathrm{d}P_{YX}}{\mathrm{d}P_{Y}P_{X}}\left(y,x\right)\mathrm{d}P_{Y}P_{X}\left(y,x\right)$
-$=$ $\int\mathds{1}_{\hat{\mathcal{A}}}(y,x)\frac{\mathrm{d}P_{YX}}{\mathrm{d}P_{Y}P_{X}}\left(y,x\right)\mathrm{d}P_{Y}P_{X}\left(y,x\right)$ (69)
-$=$ $\int\int\mathds{1}_{\hat{\mathcal{A}}_{x}}(y)\frac{\mathrm{d}P_{YX}}{\mathrm{d}P_{Y}P_{X}}\left(y,x\right)\mathrm{d}P_{Y}(y)\mathrm{d}P_{X}(x)$ (70)
-$=$ $\int\int\mathds{1}_{\hat{\mathcal{A}}_{y}}(x)\frac{\mathrm{d}P_{YX}}{\mathrm{d}P_{Y}P_{X}}\left(y,x\right)\mathrm{d}P_{X}(x)\mathrm{d}P_{Y}(y)$ (71)
-$=$ $\int\mathds{1}_{\hat{\mathcal{A}}}(x,y)\frac{\mathrm{d}P_{YX}}{\mathrm{d}P_{Y}P_{X}}\left(y,x\right)\mathrm{d}P_{X}P_{Y}(x,y)$ (72)
+$P_{XY}\left(\mathcal{A}\right)=\int_{\widehat{\mathcal{A}}}\mathrm{d}P_{YX}\left(y,x\right)$ (54)
+$=$ $\int_{\widehat{\mathcal{A}}}\frac{\mathrm{d}P_{YX}}{\mathrm{d}P_{Y}P_{X}}\left(y,x\right)\mathrm{d}P_{Y}P_{X}\left(y,x\right)$
+$=$ $\int\mathds{1}_{\widehat{\mathcal{A}}}(y,x)\frac{\mathrm{d}P_{YX}}{\mathrm{d}P_{Y}P_{X}}\left(y,x\right)\mathrm{d}P_{Y}P_{X}\left(y,x\right)$ (55)
+$=$ $\iint\mathds{1}_{\widehat{\mathcal{A}}_{x}}(y)\frac{\mathrm{d}P_{YX}}{\mathrm{d}P_{Y}P_{X}}\left(y,x\right)\mathrm{d}P_{Y}(y)\mathrm{d}P_{X}(x)$ (56)
+$=$ $\iint\mathds{1}_{\widehat{\mathcal{A}}_{y}}(x)\frac{\mathrm{d}P_{YX}}{\mathrm{d}P_{Y}P_{X}}\left(y,x\right)\mathrm{d}P_{X}(x)\mathrm{d}P_{Y}(y)$ (57)
+$=$ $\int\mathds{1}_{\widehat{\mathcal{A}}}(x,y)\frac{\mathrm{d}P_{YX}}{\mathrm{d}P_{Y}P_{X}}\left(y,x\right)\mathrm{d}P_{X}P_{Y}(x,y)$ (58)
 $=$ $\int_{\mathcal{A}}\frac{\mathrm{d}P_{YX}}{\mathrm{d}P_{Y}P_{X}}\left(y,x\right)\mathrm{d}P_{X}P_{Y}(x,y),$ (59)
-where the set $\hat{\mathcal{A}}$ is defined in (40).
-By performing a change of measure using Theorem 2 and assumption $\left(a\right)$, the equality in (68) is obtained; and the equality in (71) follows by exchanging the order of integration [5, Theorem 2.6.6].
-The proof is completed from Theorem 1 and by combining equations (56),  (61),  (67) and (73), which establish (53), (54) and (55). ∎
-Theorem 11 (Inverse Bayes-like Rule)
+where the set $\widehat{\mathcal{A}}$ is defined in (27).
+The equality in (54) is obtained by performing a change of measure using Theorem 2 under assumption $\left(a\right)$; and the equality in (57) follows by exchanging the order of integration [5, Theorem 2.6.6].
+The proof is completed from Theorem 1 and by combining equations (42),  (47),  (53) and (59), which establish (40) and (41). ∎
+Theorem 11 (Inverse Bayes-like Rule).
 Consider the conditional probability measures $P_{Y|X}$ and $P_{X|Y}$; and the probability measures $P_{Y}$ and $P_{X}$ that satisfy (31) and (32); and the joint probability measures $P_{YX}$ and $P_{XY}$ in (26) and (33) respectively.
-Assume that:
-  • $\left(a\right)$
-    For all $x\in\mathcal{X}$${P_{Y}}\ll{P_{Y|X=x}}$; and
-  • $\left(b\right)$
-    For all $y\in\mathcal{Y}$${P_{X}}\ll{P_{X|Y=y}}$.
+Assume that: $\left(a\right)$ For all $x\in\mathcal{X}$${P_{Y}}\ll{P_{Y|X=x}}$; and $\left(b\right)$ For all $y\in\mathcal{Y}$${P_{X}}\ll{P_{X|Y=y}}$.
 Then,
-$\frac{\mathrm{d}P_{X}P_{Y}}{\mathrm{d}P_{XY}}\left(x,y\right)\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{P_{XY}}\frac{\mathrm{d}P_{X}}{\mathrm{d}P_{X|Y=y}}\left(x\right)$ (74)
-=$\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{P_{XY}}$ $\frac{\mathrm{d}P_{Y}}{\mathrm{d}P_{Y|X=x}}\left(y\right)\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{P_{YX}}\frac{\mathrm{d}P_{Y}P_{X}}{\mathrm{d}P_{YX}}\left(y,x\right).$
+$\frac{\mathrm{d}P_{X}P_{Y}}{\mathrm{d}P_{XY}}\left(x,y\right)\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{P_{XY}}\frac{\mathrm{d}P_{X}}{\mathrm{d}P_{X|Y=y}}\left(x\right)\frac{\mathrm{d}P_{Y}}{\mathrm{d}P_{Y|X=x}}\left(y\right)\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{P_{YX}}\frac{\mathrm{d}P_{Y}P_{X}}{\mathrm{d}P_{YX}}\left(y,x\right).$
 Proof:
 The proof follows along the same lines as the proof of Theorem 10. ∎
-V Conclusions and Final Remarks
-This paper provides proofs for several well-known folklore theorems, as well as some other lesser-known results in information theory involving the Radon-Nikodym derivative.
-Notably, these theorems serve as fundamental tools for establishing various properties of Shannon’s information measures in a unified framework, that is, regardless of whether the underlying measures admit a probability mass function or a probability density function.
-Nonetheless, the proof of these properties are left out of the scope of this paper, due to space constraints.
-Finally, it is important to highlight that the presented results hold in full generality, as the reference measure is assumed to be any $\sigma$-finite measure.
+The following theorem introduces an identity involving the sum of a mutual information [13, 14] and a lautum information [33].
+Such a result appeared first in [42, Corollary 13] as a byproduct of the properties of the generalization error of machine learning algorithms.
+In this work, a novel proof is presented in a general context exclusively using the folklore theorems on the RND described above.
+Theorem 12.
+Consider the conditional probability measure $P_{Y|X}$; the probability measures $P_{Y}$ and $P_{X}$ that satisfy (31) and (32); and the joint probability measure $P_{YX}$ in (26).
+Let $Q$ be some $\sigma$-finite measure in $\triangle\left(\mathcal{Y}\right)$ and assume that for all $x\in\mathcal{X}$$P_{Y}\ll P_{Y|X=x}\ll Q\ll P_{Y}$.
+Then,
+$\mathrm{I}\left(P_{Y|X};P_{X}\right)+\mathrm{L}\left(P_{Y|X};P_{X}\right)$ (60)
+$=$ $\iint\log\left(\displaystyle\frac{\mathrm{d}{P_{Y|X=x}}}{\mathrm{d}{Q}}(y)\right)\mathrm{d}P_{Y|X=x}(y)\mathrm{d}P_{X}(x)$
+-- $\iint\log\left(\displaystyle\frac{\mathrm{d}{P_{Y|X=x}}}{\mathrm{d}{Q}}(y)\right)\mathrm{d}P_{Y}(y)\mathrm{d}P_{X}(x),$
+where $\mathrm{I}\left(P_{Y|X};P_{X}\right)$ and $\mathrm{L}\left(P_{Y|X};P_{X}\right)$ are respectively the mutual information and lautum information induced by the joint probability measure $P_{YX}$ in (26).
+Proof:
+From the definition of mutual information, it follows that
+$\mathrm{I}\left(P_{Y|X};P_{X}\right)=\iint\log\left(\displaystyle\frac{\mathrm{d}{P_{Y|X=x}}}{\mathrm{d}{P_{Y}}}(y)\right)\mathrm{d}P_{Y|X=x}(y)\mathrm{d}P_{X}(x)$ (61)
+$=$ $\iint\log\left(\displaystyle\frac{\mathrm{d}{P_{Y|X=x}}}{\mathrm{d}{Q}}(y)\displaystyle\frac{\mathrm{d}{Q}}{\mathrm{d}{P_{Y}}}(y)\right)\mathrm{d}P_{Y|X=x}(y)\mathrm{d}P_{X}(x)$
+$=$ $\iint\log\left(\displaystyle\frac{\mathrm{d}{P_{Y|X=x}}}{\mathrm{d}{Q}}(y)\right)\mathrm{d}P_{Y|X=x}(y)\mathrm{d}P_{X}(x)$
+$+$ $\iint\log\left(\displaystyle\frac{\mathrm{d}{Q}}{\mathrm{d}{P_{Y}}}(y)\right)\mathrm{d}P_{Y|X=x}(y)\mathrm{d}P_{X}(x),$ (63)
+where the equality in (IV) follows from Theorem 4. The proof follows by re-writing the second term in (63). That is,
+$\iint\log\left(\displaystyle\frac{\mathrm{d}{Q}}{\mathrm{d}{P_{Y}}}(y)\right)\mathrm{d}P_{Y|X=x}(y)\mathrm{d}P_{X}(x)$ (64)
+$=$ $\iint\left(\log\left(\displaystyle\frac{\mathrm{d}{Q}}{\mathrm{d}{P_{Y}}}(y)\right)\displaystyle\frac{\mathrm{d}{P_{Y|X=x}}}{\mathrm{d}{P_{Y}}}(y)\right)\mathrm{d}P_{Y}(y)\mathrm{d}P_{X}(x)$
+$=$ $\int\log\left(\displaystyle\frac{\mathrm{d}{Q}}{\mathrm{d}{P_{Y}}}(y)\right)\left(\int\displaystyle\frac{\mathrm{d}{P_{Y|X=x}}}{\mathrm{d}{P_{Y}}}(y)\mathrm{d}P_{X}(x)\right)\mathrm{d}P_{Y}(y)$ (65)
+$=$ $\int\log\left(\displaystyle\frac{\mathrm{d}{Q}}{\mathrm{d}{P_{Y}}}(y)\right)\mathrm{d}P_{Y}(y),$ (66)
+where the equality in (64) follows from Theorem 2; the equality in (65) is a reorganization of terms in the integrals; and the equality in (66) follows from Theorem 9.
+From the definition of lautum information, it follows that
+$\mathrm{L}\left(P_{Y|X};P_{X}\right)=\iint\log\left(\displaystyle\frac{\mathrm{d}{P_{Y}}}{\mathrm{d}{P_{Y|X=x}}}(y)\right)\mathrm{d}P_{Y}(y)\mathrm{d}P_{X}(x)$
+$=$ $\iint\log\left(\displaystyle\frac{\mathrm{d}{P_{Y}}}{\mathrm{d}{Q}}(y)\displaystyle\frac{\mathrm{d}{Q}}{\mathrm{d}{P_{Y|X=x}}}(y)\right)\mathrm{d}P_{Y}(y)\mathrm{d}P_{X}(x)$
+$=$ $\int\log\left(\displaystyle\frac{\mathrm{d}{P_{Y}}}{\mathrm{d}{Q}}(y)\right)\mathrm{d}P_{Y}(y)$
+$-\iint\log\left(\displaystyle\frac{\mathrm{d}{P_{Y|X=x}}}{\mathrm{d}{Q}}(y)\right)\mathrm{d}P_{Y}(y)\mathrm{d}P_{X}(x),$ (68)
+where (IV) follows from Theorem 4 and (68) is a direct consequence of the properties of the logarithmic function and Theorem 5.
+The proof is completed by plugging (66) in (63); and adding the resulting (63) to (68).
+∎
+The relevance of expressing the sum of a mutual information and a lautum information in terms of a free parameter (i.e., the measure $Q$ in (60)) lies in a hypothesis-testing interpretation for the term $\log\left(\displaystyle\frac{\mathrm{d}{P_{Y|X=x}}}{\mathrm{d}{Q}}(y)\right)$, which can be viewed as a log-likelihood ratio for distinguishing the hypothesis $Y\sim P_{Y|X=x}$ from $Y\sim Q$, given some data point $(x,y)$.
+Consequently, such a sum captures the variation of the expectation of such a log-likelihood ratio when the ground truth probability measure shifts from $P_{Y}P_{X}$ to $P_{XY}$.
+This interpretation has played a central role for establishing bridges between the generalization theory and hypothesis testing [42].